### v1

# Online Continual Graph Learning

Abstract
The aim of Continual Learning (CL) is to learn new tasks incrementally while avoiding catastrophic forgetting. Online Continual Learning (OCL) specifically focuses on learning efficiently from a continuous stream of data with shifting distribution. While recent studies explore Continual Learning on graphs exploiting Graph Neural Networks (GNNs), only few of them focus on a streaming setting. Yet, many real-world graphs evolve over time, often requiring timely and online predictions. Current approaches, however, are not well aligned with the standard OCL setting, partly due to the lack of a clear definition of online Continual Learning on graphs. In this work, we propose a general formulation for online Continual Learning on graphs, emphasizing the efficiency requirements on batch processing over the graph topology, and providing a well-defined setting for systematic model evaluation. Finally, we introduce a set of benchmarks and report the performance of several methods in the CL literature, adapted to our setting.
Index Terms:
continual learning, online learning, graph neural networkI Introduction
In standard machine learning setups, models are at first trained in an environment where all training data is simultaneously available to the learning algorithm, and only at a later time model predictions are produced on new input data. Importantly, data observations are assumed to be mutually independent and identically distributed. Real-world environments however often generate data in chunks or streams which often undergo shifts in the data distribution or even variations in tasks to be solved. In turn, previously trained models may require expensive retraining or model reconfiguration to stay up to date. In this setting, Continual Learning (CL) [1, 2], lifelong learning [3] and incremental learning [4] are similar machine learning paradigms sharing the same goal of adapting models to incrementally learn as soon as new data and new tasks are presented.
Even further, in the online learning setting, training data points are collected sequentially, processed in real-time by the learning method, and immediately discarded [5, 6]. Such strict environments are found in monitoring and control problems [7, 8], such as traffic management, activity recognition, fraud detection, with other applications such as on data generated by optical sensors [9] or prediction of power production considering environmental conditions [10]. Online CL (OCL) is therefore an extremely challenging setting that requires models to quickly adapt to new conditions, allowing for anytime inference with minimal latency and without forgetting already acquired knowledge.
Research on CL extends to graph-structured data as well, a field often referred to as Continual Graph Learning (CGL) [11]. Indeed, many machine learning tasks involve graph representations of data, such as social networks, citation networks, biological systems and transaction networks. However, most graphs in the real world are not static: they continuously evolve over time experiencing the addition/removal of nodes and changes to their topology. Examples include social network growths, new publications in citation networks, and changes in road conditions [12, 13]. However, the literature has yet to formalize a rigorous online setting for CGL, and lacks a thorough analysis of its unique challenges. Among all, node-level predictions in graph machine learning often rely on past neighboring information, as seen in Graph Neural Networks (GNNs), leading to computational challenges as graphs densify over time.
In this paper, we provide four main contributions to address this gap.
1.
We formalize the Online Continual Graph Learning (OCGL) framework, establishing a foundation for CL on graphs in a node-level streaming environment. By doing so, we bridge the gap between OCL and CGL settings.
2.
We highlight the critical issue of neighborhood expansion, which could undermine the online computational efficiency due to the complexity of multi-hop aggregation in GNNs as the graph expands. To address it, we propose and discuss a first, simple and viable solution to ensure that models can efficiently scale as the graph evolves.
3.
We present a benchmarking environment for OCGL, encompassing six datasets, and evaluate several existing methods; as most of these methods are designed for offline CL, we suitably adapt them to operate under the OCGL setting.
4.
We conduct further empirical analyses, including ablation studies considering different settings and model configurations and evaluation over heterophilic graphs. Our findings reveal higher performance of replay-based methods that are tailored to preserve topological information in the considered setting.
We believe this work lays a strong foundation for future research to develop sound and more effective approaches for online continual graph learning.
II Background and related works
Continual Learning. Depending on the type of shift in the data distribution, CL has been categorized into three main scenarios [14]: in task-incremental learning, the model sequentially learns distinct tasks, which requires availability of task identifiers to make predictions; class-incremental learning consists in classifying instances with an increasing number of classes, without task identifiers; finally, domain-incremental learning requires solving the same problem in different contexts. In the past, the main applications of CL included reinforcement learning [15, 16] and especially computer vision [17, 18, 19, 20, 21, 6], although most of the methods that have been developed to address these problem domains are not limited to these fields and can be used for a wide range of other machine learning tasks. Three main broad categories of CL approaches to mitigate forgetting have been proposed in the literature [2]: regularization methods, replay methods and architectural methods. Regularization methods [15, 22, 19, 20, 4] introduce additional loss terms to preserve important parameters to retain previously acquired knowledge. Replay methods [17, 18, 23, 5] use a memory buffer to store some representative samples from old tasks, to use them jointly with new samples to update model parameters. Architectural methods [24, 25, 26] avoid changes to model parameters either by gating mechanisms or by introducing new parameters, allowing the model to grow.
Online Continual Learning. In the usual CL scenarios described above, data arrive one task at a time, allowing for offline training with multiple passes and shuffles over the data for the current task [2]. A more realistic and challenging scenario is that of Online Continual Learning (OCL) [5, 6, 27], where data arrive in small batches of only few samples, without the possibility for the model to store all the data for the current task, either for privacy reasons or memory limitations. In this setting the learning algorithm must be able to use efficiently the mini-batches that arrive in the non-stationary stream. Additionally, whereas for CL we assume to know the task boundaries, OCL can be performed in a boundary-agnostic setting, or task-free, potentially allowing for more varied distribution shifts compared to the three settings described before [28]. However, many CL methods are not suited to this setting and require modifications. An additional characteristic of OCL is the ability to perform anytime inference: the model should always be up-to-date and ready to make predictions online after each training batch, reacting quickly to distribution shifts [28].
Learning on graphs. Graph Neural Networks (GNN) [29, 30, 31, 32] have emerged as the state-of-the-art approach for dealing with network data, generalizing convolution to graph structures. The core mechanism of most GNNs is message passing [33]: at each layer, the hidden embedding of each node is updated using information from its neighborhood as . Here AGGREGATE and UPDATE are differentiable functions specified by the particular model. Specifically, as at each step each node updates its embedding using the information (message) coming from its neighbors, after layers it will depend on its -hop neighborhood. Graph-based processing of temporal data has a relatively short history, primarily encompassing the study of temporal graphs [34, 35] and time series data [36, 37] with dedicated adaptation strategies to deal with evolving graphs [38] and benchmarks [39].
Continual Graph Learning. In the last few years, researchers have started to develop CL strategies tailored to graph data [40], with applications such as recommender systems [41] and traffic prediction [42]. The main approaches resemble general CL strategies, with a particular focus on preserving topological information with a loss term on neighborhood aggregation parameters [12], or specific node selection policies to retain informative nodes as memory buffer [13]. Recently, a number of surveys have been published on the topic [43, 11, 44, 45], and a benchmark has been proposed [46]. Continual Graph Learning (CGL) possesses some peculiarities that differentiate it from other problem domains with independent samples, as graph structure requires careful consideration. Specifically, we can distinguish between graph-level CGL and node-level CGL [46]. In the first, the model needs to make predictions about entire graphs, and thus each sample in the training data is an independent graph. Thus, standard CL settings and methods can be applied with relative ease [47]. Instead, in node-level GCL predictions are made about the single nodes in usually one single graph. In this node-level CGL setting each new task therefore consists a new subgraph of the overall graph, with new classes or type of nodes. Specifically, the task subgraph arrives all at once, and offline training is performed on it. Here, an additional specification needs to be made about the availability of inter-task edges [45]: since message passing GNNs aggregate information from neighboring nodes, the presence of inter-task edges towards nodes of previous tasks would inevitably lead to the use of past data. In practice, inter-task edges are often kept, but without access to the labels from past tasks [44]. Finally, the addition of new nodes creates an additional source of backward interference, known as structural shift [48]: when evaluating the model on nodes belonging to past tasks, their neighborhood composition and topology will be changed compared to when the task was observed.
III Online Continual Graph Learning
The OCL setting described previously has been explored in domains such as computer vision [6, 27] and sequences [49], but to the best of our knowledge, it has not yet been thoroughly studied for graph-structured data. Some papers on CGL consider a setting referred to as streaming [40, 50], yet the approaches can be categorized as offline CL as the streams consist of graph snapshots (often corresponding to entire tasks or nonetheless not classifiable as mini-batches), on which models are trained with multiple passes. While some CGL methods could in principle be used in an online fashion [43], the practical implications and constraints of the online setting for graph data have not been investigated yet. We highlight this gap in Table I and fill it with this paper.
We introduce Online Continual Graph Learning (OCGL), a new framework that ports CGL to the online problem setting. Specifically, OCGL is applicable to dynamic real-world scenarios such as social networks or recommender systems, where sudden distribution changes occur, and quick model adjustments are essential for making anytime predictions. We focus on the study of node classification, although the framework can be adapted to regression problems or edge-level tasks.
III-A Problem formulation
A growing network. We model the data associated with an OCGL problem as an evolving graph induced by a stream of nodes . At each time step , the monitored system is represented as a graph defined by node set , edge set , and a set of node attributes . Edge attributes, e.g., accounting for the strength of the connection or defining the type of node-node relations, can be considered likewise, however, they are not discussed here to ease the presentation. The graph nodes can be associated with class labels or target feature vectors collected in set to be predicted and/or used as training samples to learn the model. For simplicity, in this paper we focus on classification tasks, as most commonly done in CL. Graph is a snapshot of the temporal graph and, together with , collects all information available up to time step . At each time step a new node is acquired from the monitored system and populates the current snapshot graph . Specifically, tuple containing a new node index , associated node features , and a set of neighbors is presented and connected to graph according to the relations contained in . Finally, the target class label of node may or may not be acquired contextually to ; for instance, a prediction for node might be requested at time while the true class label is observed only at a later time.
The definition of the evolving graph is general, as it does not make assumptions on the distribution shifts happening in the node stream. This general, task-free node stream can be easily adapted or made more specific: while a real-world stream could be induced by a time-stamp on the nodes, this setting can be derived from any static graph by establishing an ordering on the nodes. The three CL scenarios of task-, class- and domain-incremental can thus be easily adapted to an online setting by ordering nodes by task, similarly to what is done in other domains [6, 27].
| Properties | CL | OCL | CGL | OCGL |
| Incremental learning | ||||
| Graph representation learning | ||||
| Single pass over stream | ||||
| Anytime predictions | ||||
| Efficiency constraints (memory) | ||||
| Efficiency constraints (computation) | ||||
| Neighborhood expansion problem |
Problem statement. The goal of OCGL is to incrementally update a prediction model , using only information associated with the current node and its neighbors of a given order within . Parameters have a predefined (maximum) dimension and model is trained incrementally as soon as new nodes and the associated labels are provided. New nodes are acquired either individually or in small mini-batches, slightly weakening the online setting as commonly done in the literature [5]. Importantly, in line with CL principles, the model must retain previously acquired knowledge: at time , the model may be asked to make predictions about , by using information from its neighbors in .
The key difference with standard CGL is that in OCGL, as in OCL, the model needs to adapt quickly in order to perform anytime predictions on the node stream [28]. This objective introduces specific efficiency constraints not present in the CGL literature; we refer the reader to Table I. The first requirement is to have small mini-batches of nodes: while in CGL settings training is done on a large task-specific subgraph, in OCGL this would be in contrast with the objective of anytime predictions. Due to their small size, mini-batches do not represent meaningful graph structures on their own. Consequently, OCGL relies on inter-batch edges. Another important requirement is that mini-batches are seen only once and, after training is performed, the mini-batch is discarded; a possible exception is the use of a fixed-size memory buffer that can store a limited number of samples. Finally, we require the training on each mini-batch to have bounded compute and memory budgets, independently from graph size at the current time step. This efficiency constraint poses non-negligible issues associated with reiterated message passing within multi-layer GNNs. We delve deeper into these complexities below, and we illustrate the difference of graph evolution compared to GCL in Figure 1.
Neighborhood expansion problem. At each of the layers of the GNN, the network aggregates the embeddings of the neighboring nodes, thus requiring access to -hop neighborhoods of nodes. The size of the -hop neighborhood of a node however scales as where is the average degree. Moreover, is time-dependent and can increase as the graph grows: real-world graphs exhibit densification over time, with the number of edges following a power law of the number of nodes, and a shrinking of graph diameter over time [51, 52]. Depending on how well-connected a graph is, very few hops may be required to go from any node to almost any other one. Thus, large or would require processing a number of nodes in the order of the entire (growing) graph for each mini-batch, going against efficiency in our definition of OCGL. It is therefore of paramount importance to keep a relatively small or to introduce a strategy to deal with a large .
To provide evidence of the neighborhood expansion phenomenon on four of the datasets considered in our experiments (see Section V for further details), in Figure 2 we plot the size of the -hop neighborhood of each mini-batch in the node stream inside the evolving graph, for some values of . The Reddit graph in particular is very well connected, with two hops containing the majority of the graph, and three hops covering practically all nodes. CoraFull and Arxiv show a more moderate neighborhood expansion, but still prohibitive. Lastly, two hops of the Amazon Computer graph covers, even in the worst case, about half of the nodes. Additional plots are shown in Appendix B, with the addition of the number of edges connecting the various hops, which can be used as a proxy of computational complexity.
Mini-batch acquisition. Within the introduced OCGL framework, and depending on the application, predictions may be requested from -hop neighbors with . However, due to the discussed neighborhood expansion problem, using all of these nodes does not allow operating within the computational efficiency constraints. This issue is similar to the problem of scaling static GNNs for large graphs, for which mini-batch training is required both for memory and efficiency reasons. Numerous approaches have been developed to address this problem, such as fixing a number of neighborhood sampled for aggregation [53, 54] or training on partitions of the graph [55]. The problem is also present in CGL, but it is not in contrast with the requirements of the setting, and is thus not discussed in the literature. In our context, the simplest solution to guarantee compliance with OCGL requirements appears to be providing a maximum number of neighboring nodes through any form of sub-sampling, thus guaranteeing an upper bound on the size of the computational graph for each batch. It follows that the OCGL model will not have direct access to the entire graph . For each node in a mini-batch, an OCGL model cannot leverage the full -layer ego-graph within . Instead, it will only received a sub-sampled .
The way in which the sub-sampled -hop neighborhoods are obtained starting from the data generating process of the growing graph may depend on the specific application, but we can generally conceptualize two distinct systems, which are nonetheless equivalent regarding the learning process of the model. In one case, an up-to-date snapshot is assumed to be stored in a Past Information Store (PIS) system – distinct from an eventual memory buffer associated with predictive model – as in the more general lifelong-learning system setup [3]. We do not impose memory limitations on the PIS to allow graph growth, but we still require the training on each mini-batch to have bounded computational time and memory cost, and assume to only have access to the labels of nodes in the current batch. In the other case, mini-batches are directly received in the form of sub-sampled neighborhoods from an external data source. This could be the case when privacy concerns are involved, or in case of huge graphs where a local exploration is performed.
IV Strategies for OCGL
Having defined the Online Continual Graph Learning setting, we consider and evaluate some popular CL techniques, most of which are agnostic with respect to the type of the input data. Some CGL learning strategies are easily adapted to the online setting, such as PDGNN [56] and SSM [57], while others are not directly applicable, such as ER-GNN [13], which stores representative nodes according to metrics computed offline on an entire graph snapshot, and thus we simplify it as described below. Similar efficiency considerations exclude many more recent CGL methods, which may require expensive steps and/or access to most of the graph: to name a few, CaT [58] requires access to the task subgraph for condensation, UGCL [59] uses a global structure distillation which requires to compute node embeddings for the entire graph, and a local structure distillation which still requires to compute the full embedding of buffer nodes and all their neighbors, and MSCGL [60] performs a costly neural architecture search for each task. Additionally, baselines that require expensive fine-tuning steps such as GDumb [61] are excluded, as they would violate the online setting. Overall, most strategies natively require task boundaries, and have been modified for the task-free setting as described below.
- 1.
2.
EWC. Elastic Weight Consolidation [15] adds a quadratic term to the loss to penalize the modification of important parameters. Parameter importance is approximated by the diagonal of the Fisher information matrix, which needs to be computed offline for each task. We therefore modify the algorithm to keep one single Fisher information matrix updated with a running average over the batches, similarly to the MAS approach detailed later. Another approach would be to keep a moving average, as done in EWC++ [4].
3.
A-GEM. Averaged GEM [5] is a more efficient version of GEM [18], which ensures that the average loss for past tasks does not increase. It achieves this by projecting the gradient of the incoming batch in the orthogonal space of the gradient computed on samples from a memory buffer, if their scalar product is negative. We select buffer nodes with reservoir sampling [62].
- 4.
5.
MAS. Memory Aware Synapses [19] is a quadratic regularization similar to EWC, but it calculates importance as the sensitivity of the output on parameters. MAS is natively an online method, as the importance scores are accumulated with each new data point.
- 6.
- 7.
8.
SSM. Sparsified Subgraph Memory [57] consists in storing sparsified computational graphs of nodes in the memory buffer, allowing to leverage topological information in replay methods. We use this memory buffer by selecting nodes via reservoir sampling [62], as support for both ER and A-GEM strategies.
V Experimental Setup
In this section we introduce the specific experimental setup used, describing the construction of the node streams from benchmark datasets, and the details of model training and evaluation.111Code available at https://github.com/giovannidonghi/OCGL.
Benchmarks. Six node classification graph datasets are used in our experiments: four homophilous multiclass datasets CoraFull [66], Arxiv [67], Reddit [53] and Amazon Computer [68], the heterophilous Roman Empire [69] and the binary classification dataset Elliptic [70]. The datasets are described in Appendix A. As is common in the OCL literature [6, 27], in order to position our experiments close to the rest of the CL literature, for all datasets except for Elliptic we devise a node stream derived from the class-incremental CL setting, which is considered as the most challenging one for catastrophic forgetting [21]. We divide the nodes in the graph into groups with fixed order consisting of 2 classes: this would be the sequence of two-class tasks in class-incremental learning (resulting in 35 tasks for CoraFull, 20 tasks for Arxiv and Reddit, 5 for Amazon Computer and 9 for Roman Empire). Then, we fix and ordering on the nodes of each task, and we stream the nodes accordingly. Therefore, the graph will gradually grow with mini-batches of nodes from two new classes at a time, which are processed in an online fashion. This allows us to consider metrics from the CL literature which require task boundaries, even though in our experiments the learning algorithm itself is task agnostic and simply adds a new output neuron when an instance of a new class is observed. For Elliptic, which only has two classes and has time information, we construct a time-incremental stream, that is according to the node real timestamps, and we divide the nodes into 10 tasks solely for the purposes of evaluation. For each dataset, we split the graph into 60% for training, 20% for validation and 20% for testing. A transductive setting is used: validation and test nodes are not used for loss computation, but they are still used for message passing.
Performance assessment. We consider three metrics widely adopted in the literature: Average Performance (AP), Average Forgetting (AF) [18], and Average Anytime Performance (AAP) [71]. AAP is obtained by evaluating the model on the validation set after each training batch, which we refer to as anytime evaluation. For the highly unbalanced Elliptic dataset, we consider the F1 score as performance metric, while for all other datasets accuracy is used. More details are reported in Appendix C.
Training details. In our experiments the backbone for all the CL strategies is the Graph Convolutional Network (GCN) [32], specifically a 2-layer GCN with a fixed hidden dimension of 256 units as done by Zhang et al. [46], with the exception of PDGNN which uses SGC [65] followed by a 2 layer MLP with 256 units. We use Adam optimizer [72] without weight decay, tuning the learning rate as a hyperparameter with the protocol defined below. We consider the batch size to be fixed, as it could depend on the real-world problem. For the smaller datasets CoraFull, Amazon Computer, and Roman Empire we consider batches of 10 nodes, while for the larger Arxiv, Reddit, and Elliptic we use 50. To evaluate the impact of batch size, in Appendix E we report experiments with larger batch sizes for each dataset. To address the neighborhood expansion problem raised in Section III-A, we perform neighborhood sampling. We fixed the number of neighbors to 10 for CoraFull, Amazon Computer, and Arxiv, and 25 for Reddit, as the latter is much more densely connected compared to the others, with an average degree of 492. As suggested by Aljundi et al. [73], multiple passes on the same mini-batch before passing to the next can be beneficial. We therefore considered as an additional tuned hyperparameter whether to perform multiple passes (5) on each mini-batch. We stress how multiple passes on an individual mini-batch are compliant with the constraints of the online setting, as they do not require storing past mini-batches beyond the current one and processing time is multiplied by a constant factor, the number of passes. As a baseline, we use a bare model that is simply fine-tuned on the incoming stream without applying any CL strategy. Additionally, we provide an upper bound in the form of a model that is jointly trained offline on the entire graph. For replay based methods, we consider different sizes of the memory buffer: for CoraFull, Amazon Computer, Roman Empire, and Elliptic we consider 1000 or 4000 nodes, and for the larger Arxiv and Reddit 1000, 4000 or 16000.
Hyperparameter selection. Many works in the CL literature use a learning protocol that is akin to the classic machine learning setting, selecting hyperparameters by performing as many full passes over the task sequence as required by a grid search. This protocol violates stricter definitions of Lifelong Learning, where the stream is observed only once, and is indeed unrealistic for real applications where a model needs to quickly adapt to changes in data distribution. Chaudhry et al. [5] therefore proposed a more sensible hyperparameter selection protocol, which has now been used in several works [41, 6, 27] and that we use for our experiments. With this protocol, only the first few tasks are used for hyperparameter selection, allowing the model to perform multiple passes, with the same online setting, over them to select the hyperparameters that lead to the best performance on validation nodes. In our case we considered approximately 20% of the tasks for this validation, with the exception of Amazon Computer where it was set to 2 as there are only 5 total tasks. Hyperparameters are then selected based on the AP on the validation set at this validation boundary (details on the search space are reported in Appendix D).
VI Results
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
In this section we discuss empirical results obtained on mini-batches with neighborhood sampling. Further results with larger batch sizes are reported in Appendix E, together with experiments with full neighborhood to assess the impact of sampling on model performance. A comparison of AP, AAP and AF for the considered CL methods on the six benchmark datasets is reported in Tables II-VII. The results for replay-based methods in the table are those obtained with the larger memory buffer sizes (4000 for CoraFull, Amazon Computer, Roman Empire, and Elliptic, 16000 for Arxiv and Reddit), while a comparison between different buffer sizes is shown in Figure 3 and in Appendix E. All experiments were repeated five times with different initializations, and the metrics are reported as mean and standard deviation across runs.
VI-A Homophilous dataset, class-incremental stream
We discuss here the results for the four homophilous datasets with a class-incremental stream, a setting that remains closer to the rest of the CGL literature [46, 57, 56].
Final Average Performance. Across all datasets, we observe that most of the compared strategies do not come close to the upper bound consisting of joint training. In general, the considered replay methods, specifically A-GEM, SSM-ER and PDGNN achieve higher final AP compared to the baseline and regularization methods, and often with a large margin. This can be expected, as rehearsal methods in CL are generally known to achieve most of the state-of-the-art results [14]; this holds true also for CGL [46]. Regularization methods struggle more than the replay strategies, as only on CoraFull EWC is significantly better than the bare baseline, while on other datasets their performances are closer to the lower bound.
Anytime Average Performance. Looking at Average Performance gives us an easy way to compare the performance after the entire learning process. However, since in the online setting we expect the model to be ready to make predictions at any time, it is arguably more useful to look at Average Anytime Performance as a metric of performance over time. For example, while on Reddit (Table IV) A-GEM records a higher final AP than PDGNN and SSM-ER, its AAP is considerably lower. Thus, in this case PDGNN and SSM-ER may be preferable, due to a more robust performance over time. For this purpose Figures 14-16 of Appendix F can give us some additional insights into the performance trends through the node stream. We note how it is natural and expected that accuracy tends to decrease with the batch index, as new classes are introduced and the classification task gets increasingly complex. In general, we observe that while the performance of regularization methods is almost monotonically decreasing, the performance of replay-based methods shows much higher variations.
Buffer size for replay methods. In Figure 3, we observe the results obtained by the replay methods with different buffer sizes. We note how SSM-ER appears to be the only method to consistently and greatly benefit from an increased buffer size. As for SSM-based buffers the number of examples is lower than the size of the buffer due to the storage of some neighboring nodes in the form of a sparsified subgraph, we can understand how with lower memory capacity this strategy may not be viable, potentially overfitting to a very small set of examples. PDGNN on the other hand shows a consistently high performance even with smaller buffer size, indicating a more efficient use of memory, as it is topology aware.
Strategy comparison. To better understand the different results in light of the choice of strategy, we report in Figure 4 a more detailed breakdown of accuracy by task for three representative methods on the Reddit dataset. With this comparison we clearly see how TWP (4a) struggles to retain past knowledge for most tasks, despite its regularization mechanism. Instead, A-GEM (4b) seems to achieve a better trade-off between stability and plasticity, maintaining the capacity to learn new tasks while generally preserving acquired knowledge. Finally, as discussed earlier PDGNN (4c) shows a more consistent performance through time, despite a lower final performance.
VI-B Heterophilous dataset, class-incremetal stream
To the best of our knowledge, the issue of heterophily in CGL has been addressed only by Zhao et al. [74]. However, their setting ignores inter-task edges, effectively making the graph less heterophilous. Since that requires unrealistic supervisory information on test nodes, we consider the Roman Empire dataset in our setting, showing results in Table VI. As can be expected, the results of all considered strategies are much lower than the upper bound. This can be the effect of two factors: first, the considered graph convolution may not be the best to deal with heteropily, and second, since nodes tend to connect with nodes of different classes, many connections are inter-task, and may thus not be observed when training on a certain class. This catastrophic drift in neighborhood composition is of course due to the class-incremental setting, which itself may not be realistic for heterophilous phenomena. While Roman Empire is not a real-world dataset, since it comes from words in a Wikipedia page, our results show that training GCNs in the OCGL setting poses significan challenges due to the drift in neighborhood composition.
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
VI-C Time-incremental stream
Compared to the abrupt and challenging drifts of the class-incremental stream, on the Elliptic dataset we have a tamer distribution shift, determined by the node timestamps, which naturally reduces the usefulness of CL. This is reflected in the results of Table VII, where many methods exhibit positive forgetting, indicating backward transfer. We also observe how in terms of final AP no CL method performs better than simple finetuning in the OCGL setting, and replay methods seem to perform worse than regularization ones, likely due to overfitting to the memory buffer. On the other hand, considering AAP and looking at Figure 18 (Appendix F), we see how PDGNN maintains an overall better and more stable performance, only with a slight drop at the end of the stream. Thus, while in this case using CL methods may not be crucial as in the more pronounced drifts, using the right strategy can still provide benefits.
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
VI-D Analysis of Design Choices
Impact of batch size. Regarding the dimensions of the node batches in the stream, we can compare the results of Tables II-VII with those in Tables XI-XV of Appendix E, which are obtained with larger batch sizes (from 10 to 50 and from 50 to 250). Overall, two considered sizes have somewhat similar results. However, we observe that replay-based methods, A-GEM in particular, tend to obtain lower performances with the larger batch sizes. For these methods, a smaller batch size could thus be a regularizing factor. The fact that very small batch sizes in an online setting can lead to relatively good performance is also encouraging for the future development of OCGL techniques.
Impact of sampling. By performing neighborhood sampling to process node mini-batches, we lose some information that might be helpful for the classification task. We expect therefore to trade some performance to remain within the efficiency constraint. Looking at Tables XVI-XIX (Appendix E) with results obtained with full neighborhood, we see that this is indeed true for most of the cases. On the Reddit dataset the performance degradation with sampling is more pronounced, as the number of neighbors was cut more substantially. There are nonetheless cases in which neighborhood sampling does not significantly affect performance, or even proves beneficial, such as on the Amazon Computer dataset.
Sensitivity to hyperparameters. In our experiments the backbone model architecture, including number of layers and hidden units, is kept fixed. We conducted an ablation study on CoraFull to assess the impact of these choices, with full results reported and discussed in Appendix G.
VII Conclusions
In this paper, we introduced the Online Continual Graph Learning (OCGL) setting, providing a formal framework that bridges the gap between the Continual Graph Learning and Online Continual Learning literature. This formulation establishes a foundation for studying graph-based learning in streaming environments where data arrives incrementally at the node level. Building on this, we raised the key issue of neighborhood expansion, proposing neighborhood sampling as a straightforward solution to bound the computational cost of training on each mini-batch even as the graph grows through time. We adapted six node classification datasets to align with the proposed OCGL framework, constructing node streams under both class-incremental and time-incremental learning scenarios. Our evaluation of suitably adapted Continual Learning methods highlights the higher performance of replay-based methods, specifically when using tailored strategies to leverage topological information to aid experience replay, yet criticalities remain in specific graph configurations. In future works, we plan to further tackle the issue of neighborhood expansion, developing tailored strategies that can ensure computational efficiency while better addressing the catastrophic forgetting problem. We further intend to consider more diverse node stream construction and additional tasks such as link prediction.
Appendix A Datasets
In the experiments for this paper, we used six node-level classification graph datasets. The CoraFull dataset [66] is a citation network where nodes represent research papers and edges indicate citations between them, with labels based on paper topics. Arxiv [67] is a larger citation network derived from arXiv papers in the Computer Science category. The Reddit dataset [53] consists of posts from different communities of the Reddit platform, where nodes represent posts, and edges connect posts commented on by the same user, forming a large interaction graph. Amazon Computer [68] is a co-purchase network, where nodes are products and edges indicate frequently co-purchased items within the computers category on Amazon. In contrast to previous dataset which are homophilous, Roman Empire [69] is an heterophilous dataset obtained from the homonymous Wikipedia page, where nodes represent words connected by syntactic relationships or adjacency in the sentence. Finally, Elliptic [70] is a Bitcoin transaction dataset, consisting of transactions connected by flows of Bitcoins. Only some of the nodes are labeled, either as licit (42,019 nodes) or illicit transaction (4,545 nodes). Summary statistics for the four graphs are reported in Table VIII.
| Dataset | CoraFull | Arxiv | Amazon Computer | Roman Empire | Elliptic | |
| # nodes | 19,793 | 169,343 | 227,853 | 13,752 | 22,662 | 203,769 |
| # edges | 130,622 | 1,166,243 | 114,615,892 | 491,722 | 32,927 | 234,355 |
| # classes | 70 | 40 | 40 | 10 | 18 | 2 |
Appendix B Neighborhood expansion
Neighborhood expansion has been identified as the main issue with the online setting for graphs. In Figure 2 of the main paper we showed the phenomenon, and we report here additional plots for all datasets and batch sizes. In particular, in Figures 6-8 we measure neighborhood expansion not only in terms of number of nodes in the -hop neighborhood, as done in Section VI of the main paper, but we additionally count the number of edges between hops, which gives us a good proxy of computational time. Interestingly, the number of edges increases more drastically than the number of nodes when incrementing the number of layers, further confirming the need for a strategy such as sampling to maintain the computational complexity bounded within the limits of the online learning setting. We do not present plots for the Roman Empire and Elliptic datasets, since their very low average degree (2.9 and 1.15 respectively) effectively eliminates the neighborhood expansion problem for a small number of layers.
Appendix C Metrics
Thanks to the construction of the node stream starting from the class-incremental setting (or the artificial definition of tasks for Elliptic), we can exploit two widely used metrics in CL: Average Performance (AP) and Average Forgetting (AF) [18]. The most comprehensive metric for CL, from which AP and AF are derived, is the performance matrix , where is the number of tasks and is the test classification performance on task after the model has observed task . AP is then defined as , and average forgetting as . AP serves as a single value to quantify the performance of the model after having observed the entire sequence of tasks, or stream in our case. AF measures the performance degradation (forgetting), that occurs from when a task was just observed to the end of training. As a performance metric, we use accuracy for all datasets except Elliptic. Since the latter is highly unbalanced, on it we use the F1 score of the illicit class.
To assess the performance of the model throughout the node stream, we also perform anytime evaluation, meaning that we evaluate the model on validation nodes after training on each mini-batch [28]. This allows us to capture the performance at any point in time, and observe also graphically how the model reacts to changes in data distribution. We measure this with Average Anytime Performance (AAP) [71], which is a generalization of average incremental accuracy for the online setting. Indicating with APt the average performance after training on batch , and having batches in total, . This can be interpreted as an Area Under the Curve accuracy score [28].
Appendix D Hyperparameters
A standard grid search was performed to select training hyperparameters for the models used in all experiments. We detail here the specific search space for each of the methods used in our comparisons. Two hyperparameters are common for all techniques: the learning rate, selected in the set , and the number of passes on each batch before passing to the next one, chosen between 1 and 5. No weight decay or dropout were used. Method specific hyperparameters are reported in Table IX, and specific details can be found in the original papers. In particular, the hyperparameters of regularization methods regulate the strenght of the regularization. For LwF a new hyperparameter has been introduced to adapt to the online setting: the number of batches after which to update the teacher model. For replay based methods we consider the proportion of memories to use with respect to each training batch, as using the entire buffer like in CGL is unfeasible in an online setting. The neighbors budget for SSM (both SSM-ER and SSM-A-GEM) refers to the number of the first- and second-hop neighbors to preserve in the memory budget, which influences the effective number of example in the buffer.
| Method | Hyperparameter candidates |
| A-GEM, ER, PDGNN | memory_proportion: {1,2,3} |
| SSM | neighbhors_budget: {(5,5), (10,10), (25,25)} |
| EWC, MAS | lambda: {} |
| LwF | lambda_dist: {0.1,1,10}; T: {0.2,2,20}, update_every: {1, 10, 100} |
| TwP | lambda_l: {}; lambda_t: {}; beta: {0.001, 0.01, 0.1} |
Appendix E Extended results
E-A Full results with larger batch size
In Tables XI-XIII, we provide extended results with the same setting configurations as the one explored in Section VI of the main paper, yet with larger mini-batch size. Compared to size 10 for CoraFull, Amazon Computer and Roman Empire, and 50 for Arxiv, Reddit and Elliptic, we report here batch-sizes 50 and 250 respectively.
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
| Method | AP% | AAPval% | AF% |
| EWC | |||
| LwF | |||
| MAS | |||
| TWP | |||
| ER | |||
| A-GEM | |||
| PDGNN | |||
| SSM-ER | |||
| SSM-A-GEM | |||
| bare | |||
| Joint | - | - |
E-B Results with full neighborhood
We report here empirical results obtained on mini-batches containing the full neighborhoods of newly presented nodes, in contrast with the main paper where neighborhood sampling was analyzed. Although the neighboring expansion problem is present, these results provide reference performance to be compared against the more realistic setting studied in the paper. For the Reddit dataset, due to the extremely high average degree, only one layer is used here for full neighborhood. Results are reported for CoraFull, Arxiv, Reddit and Amazon Computer datasets in Tables II-V, while they are not reported for Roman Empire and Elliptic as only a handful of nodes have a higher degree than 10 used for neighborhood sampling in the main experimentss of the paper. To be consistent with the use of full neighborhood information, the results reported for SSM are obtained by using the same number of seed nodes in the buffer, but with full neighborhood information as well. On the contrary, by using heavily sparsified subgraphs for the buffer in conjunction with full neighborhood for the node stream, we observed a degradation of performance, likely due to the different normalization weighting inside the GCN layer between the stream and the memory buffer.
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Joint | - | - | - | - | |||
| Method | Batch size 50 | Batch size 250 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Joint | - | - | - | - | |||
| Method | Batch size 50 | Batch size 250 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Joint | - | - | - | - | |||
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Joint | - | - | - | - | |||
E-C Impact of buffer size
For completeness of our results on the impact of buffer size on model performance, we show in Figures 10-11 plots similar to Figure 3 of the main body of the paper, showing both Average Performance and Average Anytime Performance achieved by replay methods with different buffer sizes, with both possible choices of mini-batch size as well.
Appendix F Anytime evaluation plots
We show here the plots with anytime evaluation, on all datasets and with the batch size used in the main experiments for the paper. In Figures 14-18 the lines indicate AP on validation nodes after training on each mini-batch, for all considered methods. In Figures 19-24, instead, we show the performance of each method in more detail, with a breakdown of accuracy by each task.
Appendix G Ablation study on CoraFull
In our experiments, the backbone model for all Continual Learning strategies was fixed as a 2-layer GCN with 256 hidden units, following Zhang et al. [46]. Thus, we performed a small ablation study on the CoraFull dataset to assess the impact of the number of hidden units and number of GCN layers. This is conducted only with full neighborhood aggregation.
G-A Number of hidden units
Still maintaining a 2-layer GCN, we changed the number of hidden units, from the original 256 to alternatively 128 and 512. We report the results in Tables XX-XXI. We observe how overall performance are lower than with 256 hidden units, possibly as this number may have been validated in previous works. A smaller network seems to consistently damage performance metrics, while with 512 units the comparison is more mixed.
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
G-B Number of GCN layers
We considered here a different number of GCN layers compared to our main results, 1 and 3 specifically, as on CoraFull also the 3-hop neighborhood has a relatively limited expansion. Results are reported in Tables XXII-XXIII. We observe overall lower results specifically with 3 layers, while ER and A-GEM improve using only 1 GCN layer.
Additionally, as a sort of “0-layer” baseline, we test the usage of a 2 layer MLP, equivalent to our main backbone GCN applied using the identity as adjacency matrix. This is the same as discarding graph topology information. For this reason, SSM is not used in this case. While most results are consequently greatly reduced, ER, GEM and EWC are able to maintain some performance even without using the graph topology.
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| SSM-ER | |||||||
| SSM-A-GEM | |||||||
| bare | |||||||
| Method | Batch size 10 | Batch size 50 | |||||
| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
| EWC | |||||||
| LwF | |||||||
| MAS | |||||||
| TWP | |||||||
| ER | |||||||
| A-GEM | |||||||
| bare | |||||||

### v2

@@
 Abstract
-The aim of Continual Learning (CL) is to learn new tasks incrementally while avoiding catastrophic forgetting.
-Online Continual Learning (OCL) specifically focuses on learning efficiently from a continuous stream of data with shifting distribution.
-While recent studies explore Continual Learning on graphs exploiting Graph Neural Networks (GNNs), only few of them focus on a streaming setting.
-Yet, many real-world graphs evolve over time, often requiring timely and online predictions.
-Current approaches, however, are not well aligned with the standard OCL setting, partly due to the lack of a clear definition of online Continual Learning on graphs.
-In this work, we propose a general formulation for online Continual Learning on graphs, emphasizing the efficiency requirements on batch processing over the graph topology, and providing a well-defined setting for systematic model evaluation.
-Finally, we introduce a set of benchmarks and report the performance of several methods in the CL literature, adapted to our setting.
-Index Terms:
-continual learning, online learning, graph neural networkI Introduction
-In standard machine learning setups, models are at first trained in an environment where all training data is simultaneously available to the learning algorithm, and only at a later time model predictions are produced on new input data.
-Importantly, data observations are assumed to be mutually independent and identically distributed.
-Real-world environments however often generate data in chunks or streams which often undergo shifts in the data distribution or even variations in tasks to be solved.
-In turn, previously trained models may require expensive retraining or model reconfiguration to stay up to date.
-In this setting, Continual Learning (CL) [1, 2], lifelong learning [3] and incremental learning [4] are similar machine learning paradigms sharing the same goal of adapting models to incrementally learn as soon as new data and new tasks are presented.
-Even further, in the online learning setting, training data points are collected sequentially, processed in real-time by the learning method, and immediately discarded [5, 6].
-Such strict environments are found in monitoring and control problems [7, 8], such as traffic management, activity recognition, fraud detection, with other applications such as on data generated by optical sensors [9] or prediction of power production considering environmental conditions [10].
-Online CL (OCL) is therefore an extremely challenging setting that requires models to quickly adapt to new conditions, allowing for anytime inference with minimal latency and without forgetting already acquired knowledge.
-Research on CL extends to graph-structured data as well, a field often referred to as Continual Graph Learning (CGL) [11].
-Indeed, many machine learning tasks involve graph representations of data, such as social networks, citation networks, biological systems and transaction networks.
-However, most graphs in the real world are not static: they continuously evolve over time experiencing the addition/removal of nodes and changes to their topology.
-Examples include social network growths, new publications in citation networks, and changes in road conditions [12, 13].
-However, the literature has yet to formalize a rigorous online setting for CGL, and lacks a thorough analysis of its unique challenges.
-Among all, node-level predictions in graph machine learning often rely on past neighboring information, as seen in Graph Neural Networks (GNNs), leading to computational challenges as graphs densify over time.
-In this paper, we provide four main contributions to address this gap.
+Continual Learning (CL) aims to incrementally acquire new knowledge while mitigating catastrophic forgetting.
+Within this setting, Online Continual Learning (OCL) focuses on updating models promptly and incrementally from single or small batches of observations from a data stream.
+Extending OCL to graph-structured data is crucial, as many real-world networks evolve over time and require timely, online predictions.
+However, existing continual or streaming graph learning methods typically assume access to entire graph snapshots or multiple passes over tasks, violating the efficiency constraints of the online setting.
+To address this gap, we introduce the Online Continual Graph Learning (OCGL) setting, which formalizes node-level continual learning on evolving graphs under strict memory and computational budgets.
+OCGL defines how a model incrementally processes a stream of node-level information while maintaining anytime inference and respecting resource constraints.
+We further establish a comprehensive benchmark comprising seven datasets and nine CL strategies, suitably adapted to the OCGL setting, enabling a standardized evaluation setup.
+Finally, we present a minimalistic yet competitive baseline for OCGL, inspired by our benchmarking results, that achieves strong empirical performance with high efficiency.
+I Introduction
+In standard machine learning, models are trained once on a fixed dataset, assuming independent and identically distributed samples.
+Real-world environments, however, often generate data in chunks or streams, undergoing shifts in the data distribution or even variations in tasks to be solved, which requests often impractical and expensive periodic retraining.
+Continual Learning (CL) [1, 2] addresses this challenge by enabling models to learn incrementally while retaining past knowledge and without requiring all data to remain available.
+In the more restrictive online learning setting, training data points are collected sequentially and must be processed by the learning method as soon as they appear and in real-time.
+Once processed, each sample is typically discarded leaving the learner with no access to past data [3, 4].
+Such strict environments are found in monitoring and control problems [5, 6] where decisions must be made continuously and under limited resources.
+Building on this, Online Continual Learning (OCL) represents a particularly challenging setting in which models must rapidly adapt to evolving data streams using small batches that are observed only once.
+Additionally, they are expected to support anytime inference, meaning they are expected to generate accurate predictions at arbitrary points in time, even before training can be considered completed.
+At the same time, OCL systems must operate with minimal computational and memory cost, and mitigate forgetting of previously acquired knowledge.
+Recently, CL has been extended to graph-structured data, giving rise to the field of Continual Graph Learning (CGL) [7].
+Indeed, many machine learning tasks involve graph representations of data, such as social networks, citation networks, biological systems, and transaction networks.
+Additionally, most graphs in the real world are not static: they continuously evolve, experiencing the addition/removal of nodes and changes to their topology.
+Examples include the growth of social networks, the appearance of new publications in citation networks, and changes in road conditions [8, 9].
+However, most existing CGL methods operate in offline and task-wise fashions, training on subgraph snapshots with multiple passes until convergence.
+Such settings fail to meet the core requirements of online learning – single-pass updates, limited budget, and anytime predictions – and overlook specific challenges of dynamic graphs.
+For instance, multi-hop message passing in Graph Neural Networks (GNNs) leads to unbounded computational growth as graphs densify due to the inclusion of new nodes and edges.
+We refer to this issue as the neighborhood expansion problem, which we analyze in this work.
+These additional constraints of the OCGL setting render most existing methods unsuitable for realistic, inherently evolving environments.
+To address these shortcomings, we introduce the Online Continual Graph Learning (OCGL) setting, which unifies continual and online learning principles for node-level graph streams.
+We formalize the incremental acquisition of graph information under these constraints and analyze the performance of existing methods when adapted to this setting, along with their practical implications.
+Our main contributions are summarized as follows.
 1.
-We formalize the Online Continual Graph Learning (OCGL) framework, establishing a foundation for CL on graphs in a node-level streaming environment.
-By doing so, we bridge the gap between OCL and CGL settings.
+We formalize OCGL as a principled setting bridging OCL and CGL, establishing a foundation for CL in environments characterized by data streams of node-level information.
 2.
-We highlight the critical issue of neighborhood expansion, which could undermine the online computational efficiency due to the complexity of multi-hop aggregation in GNNs as the graph expands.
-To address it, we propose and discuss a first, simple and viable solution to ensure that models can efficiently scale as the graph evolves.
+We highlight and discuss the neighborhood expansion problem, which OCGL introduces, and that can break the computation and memory requirements if not properly handled.
+We present a simple viable solution to address this problem.
 3.
-We present a benchmarking environment for OCGL, encompassing six datasets, and evaluate several existing methods; as most of these methods are designed for offline CL, we suitably adapt them to operate under the OCGL setting.
+We design a benchmarking environment for OCGL, encompassing seven datasets and several existing methods from the literature, suitably adapted to operate within the OCGL setting.
+Our findings reveal higher performance of replay-based methods that are tailored to preserve topological information.
 4.
-We conduct further empirical analyses, including ablation studies considering different settings and model configurations and evaluation over heterophilic graphs.
-Our findings reveal higher performance of replay-based methods that are tailored to preserve topological information in the considered setting.
-We believe this work lays a strong foundation for future research to develop sound and more effective approaches for online continual graph learning.
+We introduce LINEAR, a simple and lightweight method inspired by the observations drawn from the benchmark results.
+Despite its simplicity, LINEAR achieves competitive performance, and it constitutes a strong baseline to assess the quality of new OCGL methods.
+We believe that this work lays a strong foundation for systematic progress in CL, fostering the development of sound and more effective approaches for OCGL, while LINEAR offers a strong and reliable reference for future research, setting a clear benchmark that new methods should meet or exceed to demonstrate meaningful progress.
 II Background and related works
-Continual Learning.
+II-A Continual Learning
 Depending on the type of shift in the data distribution, CL has been categorized into three main scenarios [10]: in task-incremental learning, the model sequentially learns distinct tasks, which requires availability of task identifiers to make predictions; class-incremental learning consists in classifying instances with an increasing number of classes, without task identifiers; finally, domain-incremental learning requires solving the same problem in different contexts.
-In the past, the main applications of CL included reinforcement learning [15, 16] and especially computer vision [17, 18, 19, 20, 21, 6], although most of the methods that have been developed to address these problem domains are not limited to these fields and can be used for a wide range of other machine learning tasks.
-Three main broad categories of CL approaches to mitigate forgetting have been proposed in the literature [2]: regularization methods, replay methods and architectural methods.
+In the past, CL was mainly applied to reinforcement learning [11, 12] and computer vision [13, 14, 15, 16, 17, 4], but most of the methods that have been developed to address these problem domains can be used for a wide range of other machine learning tasks.
+CL approaches to mitigate forgetting fall into three general categories [2]: regularization, replay and architectural methods.
 Regularization methods [11, 18, 15, 16, 19] introduce additional loss terms to preserve important parameters to retain previously acquired knowledge.
 Replay methods [13, 14, 20, 3] use a memory buffer to store some representative samples from old tasks, to use them jointly with new samples to update model parameters.
-Architectural methods [24, 25, 26] avoid changes to model parameters either by gating mechanisms or by introducing new parameters, allowing the model to grow.
-Online Continual Learning.
-In the usual CL scenarios described above, data arrive one task at a time, allowing for offline training with multiple passes and shuffles over the data for the current task [2].
-A more realistic and challenging scenario is that of Online Continual Learning (OCL) [5, 6, 27], where data arrive in small batches of only few samples, without the possibility for the model to store all the data for the current task, either for privacy reasons or memory limitations.
-In this setting the learning algorithm must be able to use efficiently the mini-batches that arrive in the non-stationary stream.
-Additionally, whereas for CL we assume to know the task boundaries, OCL can be performed in a boundary-agnostic setting, or task-free, potentially allowing for more varied distribution shifts compared to the three settings described before [28].
+Architectural methods [21, 22, 23] avoid changes to model weights either by gating mechanisms or by introducing new parameters, allowing the model to grow.
+II-B Online Continual Learning
+In the usual CL scenarios described above, data arrive one task at a time, allowing offline training with multiple passes and shuffles over the data for the current task [2].
+Online Continual Learning (OCL) [3, 4, 24] addresses the more realistic case where data arrive in small batches of only few samples, without the possibility for the model to store all the data for the current task, either for privacy reasons or memory limitations.
+In this setting, the algorithm must efficiently learn from each mini-batch in a nonstationary stream.
+Additionally, whereas for CL we assume to know the task boundaries, OCL can be performed in a boundary-agnostic setting, or task-free, allowing for diverse distribution shifts [25].
 However, many CL methods are not suited to this setting and require modifications.
-An additional characteristic of OCL is the ability to perform anytime inference: the model should always be up-to-date and ready to make predictions online after each training batch, reacting quickly to distribution shifts [28].
-Learning on graphs.
+An additional characteristic of OCL is anytime inference: the model should always be up-to-date and ready to make predictions online after each training batch, reacting quickly to distribution shifts [25].
+II-C Learning on graphs
 Graph Neural Networks (GNN) [26, 27, 28, 29] have emerged as the state-of-the-art approach for dealing with network data, generalizing convolution to graph structures.
@@
 Specifically, as at each step each node updates its embedding using the information (message) coming from its neighbors, after layers it will depend on its -hop neighborhood.
-Graph-based processing of temporal data has a relatively short history, primarily encompassing the study of temporal graphs [34, 35] and time series data [36, 37] with dedicated adaptation strategies to deal with evolving graphs [38] and benchmarks [39].
-Continual Graph Learning.
+Graph-based processing of temporal data has a relatively short history, primarily encompassing the study of temporal graphs [31, 32, 33, 34] and time series data [35, 36] with dedicated adaptation strategies to deal with evolving graphs [37] and benchmarks [38].
+II-D Continual Graph Learning
 In the last few years, researchers have started to develop CL strategies tailored to graph data [39], with applications such as recommender systems [40] and traffic prediction [41].
-The main approaches resemble general CL strategies, with a particular focus on preserving topological information with a loss term on neighborhood aggregation parameters [12], or specific node selection policies to retain informative nodes as memory buffer [13].
+Most Continual Graph Learning (CGL) methods adapt general CL strategies, focusing on preserving topological information with a loss term on neighborhood aggregation parameters [8], or specific node selection policies to replay informative nodes [9].
 Recently, a number of surveys have been published on the topic [42, 7, 43, 44], and a benchmark has been proposed [45].
-Continual Graph Learning (CGL) possesses some peculiarities that differentiate it from other problem domains with independent samples, as graph structure requires careful consideration.
+Importantly, CGL differs from other problem domains because of dependencies introduced by graph structure, requiring careful consideration.
 Specifically, we can distinguish between graph-level CGL and node-level CGL [45].
-In the first, the model needs to make predictions about entire graphs, and thus each sample in the training data is an independent graph.
-Thus, standard CL settings and methods can be applied with relative ease [47].
-Instead, in node-level GCL predictions are made about the single nodes in usually one single graph.
-In this node-level CGL setting each new task therefore consists a new subgraph of the overall graph, with new classes or type of nodes.
+In graph-level CGL, each sample is an independent graph and standard CL methods apply directly [46], while node-level CGL performs predictions within a single evolving graph.
+In node-level CGL, each task consists of a new subgraph, for example with new classes of nodes.
 Specifically, the task subgraph arrives all at once, and offline training is performed on it.
-Here, an additional specification needs to be made about the availability of inter-task edges [45]: since message passing GNNs aggregate information from neighboring nodes, the presence of inter-task edges towards nodes of previous tasks would inevitably lead to the use of past data.
+A key issue is the treatment of inter-task edges [44]: since GNNs aggregate neighbor information, edges to nodes from previous tasks may implicitly expose past data.
 In practice, inter-task edges are often kept, but without access to the labels from past tasks [43].
-Finally, the addition of new nodes creates an additional source of backward interference, known as structural shift [48]: when evaluating the model on nodes belonging to past tasks, their neighborhood composition and topology will be changed compared to when the task was observed.
+Finally, adding new nodes changes previous neighborhoods, causing structural shift [47], a source of backward interference.
 III Online Continual Graph Learning
-The OCL setting described previously has been explored in domains such as computer vision [6, 27] and sequences [49], but to the best of our knowledge, it has not yet been thoroughly studied for graph-structured data.
-Some papers on CGL consider a setting referred to as streaming [40, 50], yet the approaches can be categorized as offline CL as the streams consist of graph snapshots (often corresponding to entire tasks or nonetheless not classifiable as mini-batches), on which models are trained with multiple passes.
-While some CGL methods could in principle be used in an online fashion [43], the practical implications and constraints of the online setting for graph data have not been investigated yet.
-We highlight this gap in Table I and fill it with this paper.
-We introduce Online Continual Graph Learning (OCGL), a new framework that ports CGL to the online problem setting.
-Specifically, OCGL is applicable to dynamic real-world scenarios such as social networks or recommender systems, where sudden distribution changes occur, and quick model adjustments are essential for making anytime predictions.
-We focus on the study of node classification, although the framework can be adapted to regression problems or edge-level tasks.
-III-A Problem formulation
-A growing network.
-We model the data associated with an OCGL problem as an evolving graph induced by a stream of nodes .
-At each time step , the monitored system is represented as a graph defined by node set , edge set , and a set of node attributes .
-Edge attributes, e.g., accounting for the strength of the connection or defining the type of node-node relations, can be considered likewise, however, they are not discussed here to ease the presentation.
-The graph nodes can be associated with class labels or target feature vectors collected in set to be predicted and/or used as training samples to learn the model.
-For simplicity, in this paper we focus on classification tasks, as most commonly done in CL.
-Graph is a snapshot of the temporal graph and, together with , collects all information available up to time step .
-At each time step a new node is acquired from the monitored system and populates the current snapshot graph .
-Specifically, tuple containing a new node index , associated node features , and a set of neighbors is presented and connected to graph according to the relations contained in .
-Finally, the target class label of node may or may not be acquired contextually to ; for instance, a prediction for node might be requested at time while the true class label is observed only at a later time.
-The definition of the evolving graph is general, as it does not make assumptions on the distribution shifts happening in the node stream.
-This general, task-free node stream can be easily adapted or made more specific: while a real-world stream could be induced by a time-stamp on the nodes, this setting can be derived from any static graph by establishing an ordering on the nodes.
-The three CL scenarios of task-, class- and domain-incremental can thus be easily adapted to an online setting by ordering nodes by task, similarly to what is done in other domains [6, 27].
+The OCL setting has been explored in domains such as computer vision [4, 24] and sequences [48], but it has not been thoroughly investigated for graph-structured data.
+Some papers on CGL consider a setting referred to as streaming [39, 49], yet the approaches can be categorized as offline CL as the streams consist of graph snapshots, on which models are trained with multiple passes.
+While some CGL methods could be used in an online fashion [42], the practical implications and constraints of the online setting for graph data have not been investigated yet.
+Table I summarizes these differences and motivates the new setting introduced here.
+We introduce Online Continual Graph Learning (OCGL), a new setting that ports CGL to the online problem setting.
+Specifically, OCGL is applicable to dynamic real-world scenarios such as social networks or recommender systems, where sudden distribution changes occur, and quick model adjustments are essential for anytime predictions.
+We focus on the study of node classification, although the setting can be adapted to regression problems or edge-level tasks.
+In this section we describe the general setting and its associated principles, requirements and challenges, while in Section V we will specify the instantiation of OCGL, such as the particular stream construction, used in our experiments.
+III-A A growing network
+We model the data associated with an OCGL problem as an evolving graph induced by a stream of nodes added in succession.
+At each time step , the graph snapshot is defined by node set , edge set , and a set of node attributes .
+Edge attributes can be considered as well, however they are not discussed here to ease the presentation.
+The graph nodes can be associated with class labels to be predicted and/or used as training samples to learn the model.
+At each time step a new node is added and connected to graph .
+Specifically, a tuple containing a new node index , associated node features , and a set of neighbors is presented and used to connect to graph according to the relations contained in .
+Finally, the target class label of node may or may not be acquired contextually to : for instance, a prediction for node might be requested at time while the true class label is observed only at a later time.
+The formulation of the evolving graph is general, as it does not make assumptions on the distribution shifts happening in the node stream.
+It can be easily adapted or made more specific: while a real-world stream could be induced by a time-stamp on the nodes, this setting can be derived from any static graph by establishing an ordering on the nodes.
+The three CL scenarios of task-, class- and domain-incremental can thus be easily adapted to this online setting by ordering nodes by task, similarly to what is done in other domains [4, 24].
 | Properties | CL | OCL | CGL | OCGL |
@@
 | Neighborhood expansion problem |
-Problem statement.
-The goal of OCGL is to incrementally update a prediction model , using only information associated with the current node and its neighbors of a given order within .
-Parameters have a predefined (maximum) dimension and model is trained incrementally as soon as new nodes and the associated labels are provided.
-New nodes are acquired either individually or in small mini-batches, slightly weakening the online setting as commonly done in the literature [5].
-Importantly, in line with CL principles, the model must retain previously acquired knowledge: at time , the model may be asked to make predictions about , by using information from its neighbors in .
+III-B Problem statement
+Given a model , the objective of OCGL is to incrementally update using only the information from the current node and its -hop neighborhood in , or using small mini-batches, slightly weakening the online setting as commonly done in the literature [3].
+In line with CL principles, while adapting to the evolving stream the model must also retain previously acquired knowledge.
 The key difference with standard CGL is that in OCGL, as in OCL, the model needs to adapt quickly in order to perform anytime predictions on the node stream [25].
-This objective introduces specific efficiency constraints not present in the CGL literature; we refer the reader to Table I.
-The first requirement is to have small mini-batches of nodes: while in CGL settings training is done on a large task-specific subgraph, in OCGL this would be in contrast with the objective of anytime predictions.
-Due to their small size, mini-batches do not represent meaningful graph structures on their own.
-Consequently, OCGL relies on inter-batch edges.
-Another important requirement is that mini-batches are seen only once and, after training is performed, the mini-batch is discarded; a possible exception is the use of a fixed-size memory buffer that can store a limited number of samples.
-Finally, we require the training on each mini-batch to have bounded compute and memory budgets, independently from graph size at the current time step.
-This efficiency constraint poses non-negligible issues associated with reiterated message passing within multi-layer GNNs.
+This objective introduces specific constraints (see also Table I):
+New nodes arrive individually or in small mini-batches that are processed once; after training, each batch is discarded, possibly except for a limited replay buffer.
+Differently from CGL, small mini-batches may not form meaningful self-contained subgraphs, hence inter-batch edges are naturally present and must be handled consistently.
+Each update must have bounded computational and memory cost, regardless of the total size of .
+The model must provide anytime inference, quickly adapting to new knowledge while retaining previously acquired one: after each update it can be queried on past nodes using the current graph state .
+Satisfying these constraints guarantees efficiency and scalability as the graph grows to arbitrary size, but it poses non-negligible issues associated with reiterated message passing within multi-layer GNNs.
 We delve deeper into these complexities below, and we illustrate the difference of graph evolution compared to GCL in Figure 1.
-Neighborhood expansion problem.
-At each of the layers of the GNN, the network aggregates the embeddings of the neighboring nodes, thus requiring access to -hop neighborhoods of nodes.
-The size of the -hop neighborhood of a node however scales as where is the average degree.
-Moreover, is time-dependent and can increase as the graph grows: real-world graphs exhibit densification over time, with the number of edges following a power law of the number of nodes, and a shrinking of graph diameter over time [51, 52].
-Depending on how well-connected a graph is, very few hops may be required to go from any node to almost any other one.
-Thus, large or would require processing a number of nodes in the order of the entire (growing) graph for each mini-batch, going against efficiency in our definition of OCGL.
-It is therefore of paramount importance to keep a relatively small or to introduce a strategy to deal with a large .
-To provide evidence of the neighborhood expansion phenomenon on four of the datasets considered in our experiments (see Section V for further details), in Figure 2 we plot the size of the -hop neighborhood of each mini-batch in the node stream inside the evolving graph, for some values of .
-The Reddit graph in particular is very well connected, with two hops containing the majority of the graph, and three hops covering practically all nodes.
-CoraFull and Arxiv show a more moderate neighborhood expansion, but still prohibitive.
-Lastly, two hops of the Amazon Computer graph covers, even in the worst case, about half of the nodes.
-Additional plots are shown in Appendix B, with the addition of the number of edges connecting the various hops, which can be used as a proxy of computational complexity.
-Mini-batch acquisition.
-Within the introduced OCGL framework, and depending on the application, predictions may be requested from -hop neighbors with .
-However, due to the discussed neighborhood expansion problem, using all of these nodes does not allow operating within the computational efficiency constraints.
-This issue is similar to the problem of scaling static GNNs for large graphs, for which mini-batch training is required both for memory and efficiency reasons.
-Numerous approaches have been developed to address this problem, such as fixing a number of neighborhood sampled for aggregation [53, 54] or training on partitions of the graph [55].
-The problem is also present in CGL, but it is not in contrast with the requirements of the setting, and is thus not discussed in the literature.
-In our context, the simplest solution to guarantee compliance with OCGL requirements appears to be providing a maximum number of neighboring nodes through any form of sub-sampling, thus guaranteeing an upper bound on the size of the computational graph for each batch.
-It follows that the OCGL model will not have direct access to the entire graph .
-For each node in a mini-batch, an OCGL model cannot leverage the full -layer ego-graph within .
-Instead, it will only received a sub-sampled .
-The way in which the sub-sampled -hop neighborhoods are obtained starting from the data generating process of the growing graph may depend on the specific application, but we can generally conceptualize two distinct systems, which are nonetheless equivalent regarding the learning process of the model.
+III-C Neighborhood expansion problem
+At each of the layers of the GNN, the network aggregates the embeddings of neighboring nodes, thus requiring access to their -hop neighborhoods.
+The neighborhood size however scales as where is the average degree, which tends to increase as the graph grows: real-world graphs exhibit densification over time, with the number of edges following a power law of the number of nodes, and a shrinking of graph diameter over time [50, 51].
+In many real graphs, only a few hops suffice to connect most nodes, so large or implies processing almost the entire graph per mini-batch, violating OCGL’s efficiency constraints.
+It is therefore of utmost importance to limit or to introduce a strategy to deal with .
+We note that neighborhood expansion is also present in CGL, but it is not in contrast with the requirements of that setting, and is thus not discussed in the literature.
+To illustrate this phenomenon, Figure 2 shows the size of the -hop neighborhood across mini-batches in the node stream for four datasets considered in our experiments (see Section V for further details).
+The Reddit graph in particular is very well connected, with two hops containing the majority of the graph.
+CoraFull and Arxiv show a more moderate neighborhood expansion, while two hops on the Amazon Computer graph cover about half of the nodes.
+III-D Mini-batch acquisition
+As discussed, the growth of node neighborhoods makes it infeasible to aggregate information from all -hop neighbors while maintaining a bounded computation.
+To satisfy the efficiency requirements, the model must therefore operate on a limited subset of nodes.
+This issue is similar to the problem of scaling static GNNs for large graphs, where mini-batch training is required both for memory and efficiency reasons.
+Numerous approaches have been developed, such as fixing a number of neighbors sampled for aggregation [52, 53] or training on partitions of the graph [54].
+In our context, the simplest solution to guarantee compliance with OCGL requirements is to cap the number of neighboring nodes through any form of sub-sampling, thus guaranteeing an upper bound on the size of the computational graph for each batch.
+Thus, for any node , an OCGL model cannot leverage the full -layer ego-graph within .
+Instead, it will only receive a sub-sampled .
+The way in which the sub-sampled -hop neighborhoods are obtained starting from the data generating process of the growing graph may depend on the specific application, but we can generally conceptualize two distinct systems, which are nevertheless equivalent regarding the learning process of the model.
 In one case, an up-to-date snapshot is assumed to be stored in a Past Information Store (PIS) system – distinct from an eventual memory buffer associated with predictive model – as in the more general lifelong-learning system setup [55].
@@
 IV Strategies for OCGL
+IV-A Adaptation of existing methods
 Having defined the Online Continual Graph Learning setting, we consider and evaluate some popular CL techniques, most of which are agnostic with respect to the type of the input data.
@@
 Additionally, baselines that require expensive fine-tuning steps such as GDumb [61] are excluded, as they would violate the online setting.
-Overall, most strategies natively require task boundaries, and have been modified for the task-free setting as described below.
-- 1.
-2.
+Several strategies natively require task boundaries, and have been modified for the task-free setting as described below.
+A-GEM.
+Averaged GEM [3] is a more efficient version of GEM [14], which ensures that the average loss for past tasks does not increase.
+It achieves this by projecting the gradient of the incoming batch in the orthogonal space of the gradient computed on samples from a memory buffer, if their scalar product is negative.
+We select buffer nodes with reservoir sampling [62]
 EWC.
@@
 Another approach would be to keep a moving average, as done in EWC++ [19].
-3.
-A-GEM.
-Averaged GEM [5] is a more efficient version of GEM [18], which ensures that the average loss for past tasks does not increase.
-It achieves this by projecting the gradient of the incoming batch in the orthogonal space of the gradient computed on samples from a memory buffer, if their scalar product is negative.
-We select buffer nodes with reservoir sampling [62].
-- 4.
-5.
 MAS.
@@
 MAS is natively an online method, as the importance scores are accumulated with each new data point.
-- 6.
-- 7.
-8.
 SSM.
@@
 We use this memory buffer by selecting nodes via reservoir sampling [62], as support for both ER and A-GEM strategies.
+IV-B A new baseline method
+The main purpose of this section is introducing a minimal, efficient and competitive method that can serves as reference for future OCGL research.
+Following the insights obtained from our results in Section VI, we introduce LINEAR (Lightweight Incremental NEighborhood Aggregation with Replay), a new baseline inspired by PDGNN and pushing its design philosophy to the limit with regard to simplicity.
+For each node, LINEAR first computes a feature vector by averaging features of its neighbors.
+A lightweight linear classifier is then applied to these aggregated features.
+More specifically, LINEAR consists of a single SGC layer [64] that averages up to feature vectors sampled from the 1-hop neighborhoods, followed by a multinomial logistic regression classifier – no hidden layers, no nonlinearities, and no iterative message passing are employed; is a model hyperparameter to meet the requested computational budget as described in Section III-D.
+We couple this with an experience replay mechanism, storing directly the neighborhood-averaged features via reservoir sampling.
+The computational complexity of LINEAR is significantly lower than standard GNN-based models.
+Assume for convenience fixed feature dimension and number of sampled neighbors throughout all layers, and a -class linear classifier on top.
+Due to message passing, an -layer GNN requires processing nodes to compute the embedding of a single target node.
+Each node transformation applies an linear map, and when this cost dominates both aggregation and classification.
+The overall complexity of predicting the class of a single node is therefore , reflecting the neighborhood expansion problem.
+In contrast, LINEAR requires only cost, coming from the neighbor aggregation and the final linear classifier, respectively.
+Despite its simplicity, we show in Section VI that LINEAR exhibits strong empirical performance in the OCGL benchmarks, often surpassing GNN-based methods.
 V Experimental Setup
-In this section we introduce the specific experimental setup used, describing the construction of the node streams from benchmark datasets, and the details of model training and evaluation.111Code available at https://github.com/giovannidonghi/OCGL.
-Benchmarks.
-Six node classification graph datasets are used in our experiments: four homophilous multiclass datasets CoraFull [66], Arxiv [67], Reddit [53] and Amazon Computer [68], the heterophilous Roman Empire [69] and the binary classification dataset Elliptic [70].
+In this section, we introduce the specific experimental setup used, describing the construction of the node streams from benchmark datasets, and the details of model training and evaluation.
+V-A Benchmarks
+Seven node classification graph datasets are used in our experiments: five homophilous multiclass datasets CoraFull [66], Arxiv [67], Reddit [52], Amazon Computer [68] and Products [67], the heterophilous Roman Empire [69] and the binary classification dataset Elliptic [70].
 The datasets are described in Appendix A.
-As is common in the OCL literature [6, 27], in order to position our experiments close to the rest of the CL literature, for all datasets except for Elliptic we devise a node stream derived from the class-incremental CL setting, which is considered as the most challenging one for catastrophic forgetting [21].
-We divide the nodes in the graph into groups with fixed order consisting of 2 classes: this would be the sequence of two-class tasks in class-incremental learning (resulting in 35 tasks for CoraFull, 20 tasks for Arxiv and Reddit, 5 for Amazon Computer and 9 for Roman Empire).
-Then, we fix and ordering on the nodes of each task, and we stream the nodes accordingly.
+As is common in the OCL literature [4, 24], in order to position our experiments close to the rest of the CL literature, for all datasets except for Elliptic we devise a node stream derived from the class-incremental CL setting, which is considered the most challenging one for catastrophic forgetting [17].
+We divide the nodes in the graph into groups with fixed order consisting of 2 classes: this would be the sequence of two-class tasks in class-incremental learning (resulting in 35 tasks for CoraFull, 20 tasks for Arxiv and Reddit, 5 for Amazon Computer, 23 for Products and 9 for Roman Empire).
+Then, we fix an ordering on the nodes of each task, and we stream the nodes accordingly.
 Therefore, the graph will gradually grow with mini-batches of nodes from two new classes at a time, which are processed in an online fashion.
 This allows us to consider metrics from the CL literature which require task boundaries, even though in our experiments the learning algorithm itself is task agnostic and simply adds a new output neuron when an instance of a new class is observed.
-For Elliptic, which only has two classes and has time information, we construct a time-incremental stream, that is according to the node real timestamps, and we divide the nodes into 10 tasks solely for the purposes of evaluation.
+For Elliptic and Arxiv, that have time information available, we construct a time-incremental stream, that is according to the node real timestamps, and we divide the nodes into 10 tasks solely for the purposes of evaluation.
 For each dataset, we split the graph into 60% for training, 20% for validation and 20% for testing.
 A transductive setting is used: validation and test nodes are not used for loss computation, but they are still used for message passing.
-Performance assessment.
-We consider three metrics widely adopted in the literature: Average Performance (AP), Average Forgetting (AF) [18], and Average Anytime Performance (AAP) [71].
-AAP is obtained by evaluating the model on the validation set after each training batch, which we refer to as anytime evaluation.
+V-B Performance assessment
+We consider three widely adopted metrics in the literature: Average Performance (AP), Average Forgetting (AF) [14], and Average Anytime Performance (AAP) [71].
+AAP is obtained by evaluating the model after each training batch, which we refer to as anytime evaluation, and thus allows us to understand model performance over time.
 For the highly unbalanced Elliptic dataset, we consider the F1 score as performance metric, while for all other datasets accuracy is used.
-More details are reported in Appendix C.
-Training details.
-In our experiments the backbone for all the CL strategies is the Graph Convolutional Network (GCN) [32], specifically a 2-layer GCN with a fixed hidden dimension of 256 units as done by Zhang et al.
-[46], with the exception of PDGNN which uses SGC [65] followed by a 2 layer MLP with 256 units.
+More details are reported in Appendix B.
+V-C Training details
+In our experiments the backbone for all CL strategies is the Graph Convolutional Network (GCN) [29], specifically a 2-layer GCN with a fixed hidden dimension of 256 units as done by Zhang et al.
+[45], with the exception of PDGNN which uses 2 layers of SGC [64] followed by a 2 layer MLP with 256 units, and LINEAR which simply averages features in the immediate neighborhood and uses a linear classifier on top.
 We use Adam optimizer [72] without weight decay, tuning the learning rate as a hyperparameter with the protocol defined below.
 We consider the batch size to be fixed, as it could depend on the real-world problem.
-For the smaller datasets CoraFull, Amazon Computer, and Roman Empire we consider batches of 10 nodes, while for the larger Arxiv, Reddit, and Elliptic we use 50.
-To evaluate the impact of batch size, in Appendix E we report experiments with larger batch sizes for each dataset.
-To address the neighborhood expansion problem raised in Section III-A, we perform neighborhood sampling.
-We fixed the number of neighbors to 10 for CoraFull, Amazon Computer, and Arxiv, and 25 for Reddit, as the latter is much more densely connected compared to the others, with an average degree of 492.
+For the smaller datasets CoraFull, Amazon Computer, and Roman Empire we consider batches of 10 nodes, while for the larger Arxiv, Reddit, and Elliptic we use 50, and for the much larger Products we use 250 nodes.
+To address the neighborhood expansion problem raised in Section III, we perform neighborhood sampling, fixing the number of neighbors to 10.
 As suggested by Aljundi et al.
@@
 Additionally, we provide an upper bound in the form of a model that is jointly trained offline on the entire graph.
-For replay based methods, we consider different sizes of the memory buffer: for CoraFull, Amazon Computer, Roman Empire, and Elliptic we consider 1000 or 4000 nodes, and for the larger Arxiv and Reddit 1000, 4000 or 16000.
-Hyperparameter selection.
+For replay based methods, we consider different sizes of the memory buffer: 1%, 2% and 4% of the nodes, except for the much larger Products where we use 0.1%, 0.2% and 0.4%, to avoid having an unrealistically large buffer of several tens of thousands of nodes.
+V-D Hyperparameter selection
 Many works in the CL literature use a learning protocol that is akin to the classic machine learning setting, selecting hyperparameters by performing as many full passes over the task sequence as required by a grid search.
@@
 With this protocol, only the first few tasks are used for hyperparameter selection, allowing the model to perform multiple passes, with the same online setting, over them to select the hyperparameters that lead to the best performance on validation nodes.
-In our case we considered approximately 20% of the tasks for this validation, with the exception of Amazon Computer where it was set to 2 as there are only 5 total tasks.
-Hyperparameters are then selected based on the AP on the validation set at this validation boundary (details on the search space are reported in Appendix D).
+In our case, we considered approximately 20% of the tasks for this validation, with the exception of Amazon Computer where it was set to 2 as there are only 5 total tasks.
+Hyperparameters are then selected based on AP on the validation set at this validation boundary (details on the hyperparameters are reported in Appendix C).
 VI Results
-| Method | AP% | AAPval% | AF% |
+| Method | AAP% | AP% | AF% |
+| A-GEM | |||
+| ER | |||
 | EWC | |||
@@
 | MAS | |||
+| PDGNN | |||
+| SSM-A-GEM | |||
+| SSM-ER | |||
 | TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
+| LINEAR | |||
 | bare | |||
 | Joint | - | - |
-| Method | AP% | AAPval% | AF% |
+| Method | AAP% | AP% | AF% |
+| A-GEM | |||
+| ER | |||
 | EWC | |||
@@
 | MAS | |||
+| PDGNN | |||
+| SSM-A-GEM | |||
+| SSM-ER | |||
 | TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
+| LINEAR | |||
 | bare | |||
 | Joint | - | - |
-| Method | AP% | AAPval% | AF% |
+| Method | AAP% | AP% | AF% |
+| A-GEM | |||
+| ER | |||
 | EWC | |||
@@
 | MAS | |||
+| PDGNN | |||
+| SSM-A-GEM | |||
+| SSM-ER | |||
 | TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
+| LINEAR | |||
 | bare | |||
 | Joint | - | - |
-| Method | AP% | AAPval% | AF% |
+| Method | AAP% | AP% | AF% |
+| A-GEM | |||
+| ER | |||
 | EWC | |||
@@
 | MAS | |||
+| PDGNN | |||
+| SSM-A-GEM | |||
+| SSM-ER | |||
 | TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
+| LINEAR | |||
 | bare | |||
 | Joint | - | - |
-In this section we discuss empirical results obtained on mini-batches with neighborhood sampling.
-Further results with larger batch sizes are reported in Appendix E, together with experiments with full neighborhood to assess the impact of sampling on model performance.
-A comparison of AP, AAP and AF for the considered CL methods on the six benchmark datasets is reported in Tables II-VII.
-The results for replay-based methods in the table are those obtained with the larger memory buffer sizes (4000 for CoraFull, Amazon Computer, Roman Empire, and Elliptic, 16000 for Arxiv and Reddit), while a comparison between different buffer sizes is shown in Figure 3 and in Appendix E.
+| Method | AAP% | AP% | AF% |
+| A-GEM | |||
+| ER | |||
+| EWC | |||
+| LwF | |||
+| MAS | |||
+| PDGNN | |||
+| SSM-A-GEM | |||
+| SSM-ER | |||
+| TWP | |||
+| LINEAR | |||
+| bare | |||
+| Joint | - | - |
+In this section, we discuss empirical results of the experiments described in Section V.
+A comparison of AAP, AP and AF for the considered CL methods on the various benchmarks is reported in Tables II-VIII.
+The results for replay-based methods in the tables are those obtained with the larger memory buffer sizes (4% of nodes, or 0.4% on Products), while a comparison between different buffer sizes is shown in Figure 5 and in Appendix E.
 All experiments were repeated five times with different initializations, and the metrics are reported as mean and standard deviation across runs.
-VI-A Homophilous dataset, class-incremental stream
-We discuss here the results for the four homophilous datasets with a class-incremental stream, a setting that remains closer to the rest of the CGL literature [46, 57, 56].
-Final Average Performance.
-Across all datasets, we observe that most of the compared strategies do not come close to the upper bound consisting of joint training.
-In general, the considered replay methods, specifically A-GEM, SSM-ER and PDGNN achieve higher final AP compared to the baseline and regularization methods, and often with a large margin.
+In the tables, the best results are highlighted in bold, the second-best underlined, and the third with dashed underline.
+VI-A Overall results
+Across benchmarks, the results confirm the difficulty of the OCGL setting, especially for a class-incremental stream: none of the considered strategies approaches the upper bound of AP provided by joint offline training, except for the proposed LINEAR on Amazon Computer.
+In general, the considered replay methods, specifically SSM-ER, PDGNN and LINEAR achieve higher performance compared to the baseline and regularization methods, and often with a large margin.
 This can be expected, as rehearsal methods in CL are generally known to achieve most of the state-of-the-art results [10]; this holds true also for CGL [45].
-Regularization methods struggle more than the replay strategies, as only on CoraFull EWC is significantly better than the bare baseline, while on other datasets their performances are closer to the lower bound.
-Anytime Average Performance.
+Regularization methods struggle more than the replay strategies, as only in few cases, such as with MAS on CoraFull, they score significantly better than the bare baseline, while their performances are often closer to the lower bound of simple fine-tuning.
+VI-B Anytime evaluation
 Looking at Average Performance gives us an easy way to compare the performance after the entire learning process.
-However, since in the online setting we expect the model to be ready to make predictions at any time, it is arguably more useful to look at Average Anytime Performance as a metric of performance over time.
-For example, while on Reddit (Table IV) A-GEM records a higher final AP than PDGNN and SSM-ER, its AAP is considerably lower.
-Thus, in this case PDGNN and SSM-ER may be preferable, due to a more robust performance over time.
-For this purpose Figures 14-16 of Appendix F can give us some additional insights into the performance trends through the node stream.
-We note how it is natural and expected that accuracy tends to decrease with the batch index, as new classes are introduced and the classification task gets increasingly complex.
-In general, we observe that while the performance of regularization methods is almost monotonically decreasing, the performance of replay-based methods shows much higher variations.
-Buffer size for replay methods.
-In Figure 3, we observe the results obtained by the replay methods with different buffer sizes.
-We note how SSM-ER appears to be the only method to consistently and greatly benefit from an increased buffer size.
-As for SSM-based buffers the number of examples is lower than the size of the buffer due to the storage of some neighboring nodes in the form of a sparsified subgraph, we can understand how with lower memory capacity this strategy may not be viable, potentially overfitting to a very small set of examples.
-PDGNN on the other hand shows a consistently high performance even with smaller buffer size, indicating a more efficient use of memory, as it is topology aware.
-Strategy comparison.
-To better understand the different results in light of the choice of strategy, we report in Figure 4 a more detailed breakdown of accuracy by task for three representative methods on the Reddit dataset.
-With this comparison we clearly see how TWP (4a) struggles to retain past knowledge for most tasks, despite its regularization mechanism.
-Instead, A-GEM (4b) seems to achieve a better trade-off between stability and plasticity, maintaining the capacity to learn new tasks while generally preserving acquired knowledge.
-Finally, as discussed earlier PDGNN (4c) shows a more consistent performance through time, despite a lower final performance.
-VI-B Heterophilous dataset, class-incremetal stream
+However, since in the online setting we expect the model to be ready to make predictions at any time, Average Anytime Performance is a more useful metric of performance over time.
+Although overall AP and AAP rankings agree, there are cases in which they differ.
+Figure 3 can give us some additional insights into the performance trends through the node stream.
+For example, on CoraFull we have that LINEAR scores a lower final AP compared to PDGNN, despite having a higher AAP.
+Yet, looking at Figure 3a we see that this is due to a drop in performance just at the end of the stream, while for the rest of the time LINEAR stays consistently above PDGNN.
+Or in Figure 3b, for the Reddit dataset, we see how while LINEAR overall closely follows PDGNN, it suffers from significantly more contained drops at task boundaries.
+Thus, these cases confirm AAP as more suitable metric than AP, as the behavior of LINEAR is preferable for anytime evaluation.
+VI-C Strategy comparison
+To better understand the different results in light of the choice of strategy, we report in Figure 4 a more detailed breakdown of accuracy by task for three representative methods on the CoraFull dataset.
+With this comparison we clearly see how MAS (4a) retains past knowledge for some tasks thanks to its regularization, but struggles to learn new information.
+Instead, PDGNN (4b) and LINEAR (4c) strike a better balance between stability and plasticity.
+They obtain similar results, but as observed above LINEAR appears more robust to sudden performance drops compared to PDGNN.
+VI-D Time-incremental stream
+Compared to the abrupt and challenging drifts of the class-incremental stream, with a time-incremental stream, determined by the node timestamps, we have a tamer distribution shift which naturally reduces the usefulness of CL.
+This is reflected in the results of Tables VII-VIII, where many methods exhibit positive forgetting, indicating some backward transfer.
+Although CL offers limited improvement in these conditions, considering AAP and looking at appendix Figure 12 (Appendix E), we see how PDGNN maintains an overall better and more stable performance, only with a slight drop at the end of the stream.
+Thus, while on these two dataset using CL methods may not be crucial as in cases with more pronounced drifts, using the right strategy can still provide benefits.
+| Method | AAP% | AP% | AF% |
+| A-GEM | |||
+| ER | |||
+| EWC | |||
+| LwF | |||
+| MAS | |||
+| PDGNN | |||
+| SSM-A-GEM | |||
+| SSM-ER | |||
+| TWP | |||
+| LINEAR | |||
+| bare | |||
+| Joint | - | - |
+| Method | AAP% | AP% | AF% |
+| A-GEM | |||
+| ER | |||
+| EWC | |||
+| LwF | |||
+| MAS | |||
+| PDGNN | |||
+| SSM-A-GEM | |||
+| SSM-ER | |||
+| TWP | |||
+| LINEAR | |||
+| bare | |||
+| Joint | - | - |
+VI-E Heterophily
 To the best of our knowledge, the issue of heterophily in CGL has been addressed only by Zhao et al.
@@
 However, their setting ignores inter-task edges, effectively making the graph less heterophilous.
-Since that requires unrealistic supervisory information on test nodes, we consider the Roman Empire dataset in our setting, showing results in Table VI.
+Since that requires unrealistic supervisory information on test nodes, we consider the Roman Empire dataset in our setting, showing results in Table IX.
 As can be expected, the results of all considered strategies are much lower than the upper bound.
-This can be the effect of two factors: first, the considered graph convolution may not be the best to deal with heteropily, and second, since nodes tend to connect with nodes of different classes, many connections are inter-task, and may thus not be observed when training on a certain class.
+This can be the effect of two factors: first, the considered graph convolution may not be the best to deal with heterophily, and second, since nodes tend to connect with nodes of different classes, many connections are inter-task, and may thus not be observed when training on a certain class.
 This catastrophic drift in neighborhood composition is of course due to the class-incremental setting, which itself may not be realistic for heterophilous phenomena.
-While Roman Empire is not a real-world dataset, since it comes from words in a Wikipedia page, our results show that training GCNs in the OCGL setting poses significan challenges due to the drift in neighborhood composition.
-| Method | AP% | AAPval% | AF% |
+While Roman Empire is not a real-world dataset, since it comes from words in a Wikipedia page, our results show that training GCNs in the OCGL setting poses significant challenges due to the drift in neighborhood composition.
+| Method | AAP% | AP% | AF% |
+| A-GEM | |||
+| ER | |||
 | EWC | |||
@@
 | MAS | |||
+| PDGNN | |||
+| SSM-A-GEM | |||
+| SSM-ER | |||
 | TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
+| LINEAR | |||
 | bare | |||
 | Joint | - | - |
-VI-C Time-incremental stream
-Compared to the abrupt and challenging drifts of the class-incremental stream, on the Elliptic dataset we have a tamer distribution shift, determined by the node timestamps, which naturally reduces the usefulness of CL.
-This is reflected in the results of Table VII, where many methods exhibit positive forgetting, indicating backward transfer.
-We also observe how in terms of final AP no CL method performs better than simple finetuning in the OCGL setting, and replay methods seem to perform worse than regularization ones, likely due to overfitting to the memory buffer.
-On the other hand, considering AAP and looking at Figure 18 (Appendix F), we see how PDGNN maintains an overall better and more stable performance, only with a slight drop at the end of the stream.
-Thus, while in this case using CL methods may not be crucial as in the more pronounced drifts, using the right strategy can still provide benefits.
-| Method | AP% | AAPval% | AF% |
-| EWC | |||
-| LwF | |||
-| MAS | |||
-| TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
-| bare | |||
-| Joint | - | - |
-VI-D Analysis of Design Choices
+VI-F LINEAR
+Despite its simplicity, LINEAR reliably obtains very good results: it ranks first in four out of eight benchmarks and within the top three in all but one.
+The only exception is Arxiv with time-incremental stream, where LINEAR underperforms compared to other methods.
+In this case the distribution drift is limited, as can be inferred by the positive forgetting in Table VIII and observed in appendix Figures 14 and 21, so even the bare 2-layer GCN baseline without CL strategy performs well.
+Instead, LINEAR here suffers from its limited expressive power, since it only aggregates 1-hop information, and with a linear classifier.
+Except for this limit case, we believe that such simplicity is precisely what allows LINEAR to overperform elsewhere: it makes it more robust to forgetting, and faster to adapt.
+Additionally, its simplicity also results in much lower computational requirements, making it a suitable choice when compute is limited or a high update frequency is required.
+VI-G Design choices and ablation results
+In the previous subsections, we have discussed empirical results obtained on mini-batches with neighborhood sampling and with some fixed design choices.
+Ablation results with respect to some of these choices are provided in Appendix D, on the datasets CoraFull, Amazon Computer and Arxiv.
+We briefly discuss here the impact of these configuration decisions.
+Memory buffer size.
+In Figure 5, we observe the results obtained by the replay methods with different buffer sizes.
+While overall a larger memory buffer is associated with better performance, this is significantly visible only on CoraFull, as it is a smaller dataset with a large number of classes, and therefore a larger number of samples is required to be representative enough.
+The need for an increased buffer size is apparent for SSM-based buffers, especially for SSM-ER: as the number of examples is lower than the size of the buffer due to the storage of some neighboring nodes in the form of a sparsified subgraph, we can understand how with lower memory capacity this strategy may not be viable, potentially overfitting to a very small set of examples.
+PDGNN and LINEAR on the other hand show a consistently high performance even with smaller buffer size, indicating a more efficient use of memory, as it is topology aware.
 Impact of batch size.
-Regarding the dimensions of the node batches in the stream, we can compare the results of Tables II-VII with those in Tables XI-XV of Appendix E, which are obtained with larger batch sizes (from 10 to 50 and from 50 to 250).
-Overall, two considered sizes have somewhat similar results.
-However, we observe that replay-based methods, A-GEM in particular, tend to obtain lower performances with the larger batch sizes.
-For these methods, a smaller batch size could thus be a regularizing factor.
-The fact that very small batch sizes in an online setting can lead to relatively good performance is also encouraging for the future development of OCGL techniques.
+Regarding the dimensions of the node batches in the stream, we can compare the results of Tables II, V and III with those in Table XIII of Appendix D, which are obtained with larger batch sizes (from 10 to 50 and from 50 to 250).
+Overall, the two considered sizes have limited differences, as only on CoraFull we observe generally better performances with the larger batch size.
+On this dataset we also see a significant increase in the performance of regularization methods, with MAS becoming more competitive with the larger batch size.
+Nonetheless, the fact that very small batch sizes in an online setting can lead to relatively good performance is also encouraging for the future development of OCGL techniques.
 Impact of sampling.
@@
 We expect therefore to trade some performance to remain within the efficiency constraint.
-Looking at Tables XVI-XIX (Appendix E) with results obtained with full neighborhood, we see that this is indeed true for most of the cases.
-On the Reddit dataset the performance degradation with sampling is more pronounced, as the number of neighbors was cut more substantially.
-There are nonetheless cases in which neighborhood sampling does not significantly affect performance, or even proves beneficial, such as on the Amazon Computer dataset.
+Looking at appendix Table XIV (Appendix D) with results obtained with full neighborhood, we see that this is indeed generally true, although to a very limited extent.
+There are actually cases in which neighborhood sampling does not significantly affect performance, or even proves beneficial, possibly acting as a regularizer.
 Sensitivity to hyperparameters.
 In our experiments the backbone model architecture, including number of layers and hidden units, is kept fixed.
-We conducted an ablation study on CoraFull to assess the impact of these choices, with full results reported and discussed in Appendix G.
+We conducted an ablation study to assess the impact of these choices, evaluating also a backbone GCN with 1 or 3 layers, or with 128 and 512 hidden units, with full results reported in Appendix D.
+While changing number of hidden units does not lead to particularly relevant changes, the use of 3 layers of GCN instead of 2 generally reduces performance, in agreement with literature suggesting that deeper networks are more prone to forgetting [75].
+Results with 1 GCN layer are mixed, but still lower than the ones for LINEAR.
 VII Conclusions
-In this paper, we introduced the Online Continual Graph Learning (OCGL) setting, providing a formal framework that bridges the gap between the Continual Graph Learning and Online Continual Learning literature.
-This formulation establishes a foundation for studying graph-based learning in streaming environments where data arrives incrementally at the node level.
-Building on this, we raised the key issue of neighborhood expansion, proposing neighborhood sampling as a straightforward solution to bound the computational cost of training on each mini-batch even as the graph grows through time.
-We adapted six node classification datasets to align with the proposed OCGL framework, constructing node streams under both class-incremental and time-incremental learning scenarios.
-Our evaluation of suitably adapted Continual Learning methods highlights the higher performance of replay-based methods, specifically when using tailored strategies to leverage topological information to aid experience replay, yet criticalities remain in specific graph configurations.
-In future works, we plan to further tackle the issue of neighborhood expansion, developing tailored strategies that can ensure computational efficiency while better addressing the catastrophic forgetting problem.
-We further intend to consider more diverse node stream construction and additional tasks such as link prediction.
+In this paper, we introduced the Online Continual Graph Learning (OCGL) setting, which bridges the gap between Continual Graph Learning and Online Continual Learning.
+The OCGL formulation establishes a foundation for studying graph-based learning in streaming environments where data arrives sequentially, providing information at the level of nodes.
+A key problem that emerges in this setting is that of neighborhood expansion, which we addressed with neighborhood sampling as a straightforward solution to bound the computational and memory cost of training on each mini-batch, even as the graph grows through time.
+To support research in this area, we developed a benchmarking environment that adapted seven node-classification datasets to align with the proposed OCGL setting, constructing node streams under both class-incremental and time-incremental learning scenarios.
+Our evaluation compares nine suitably adapted methods from the CL literature along with LINEAR, a newly proposed simple and efficient baseline.
+The results indicate that replay-based methods perform best overall, especially when using tailored strategies to capture topological information in the memory buffer, while LINEAR achieves highly competitive results despite its simplicity, establishing it as a strong reference point for forthcoming research.
+In future works, we plan to further study the neighborhood expansion problem, developing tailored strategies that can ensure computational efficiency while better addressing catastrophic forgetting.
+We also intend to consider more diverse node stream construction and additional tasks such as link prediction.
 Appendix A Datasets
-In the experiments for this paper, we used six node-level classification graph datasets.
+In the experiments for this paper, we used seven node-level classification graph datasets.
 The CoraFull dataset [66] is a citation network where nodes represent research papers and edges indicate citations between them, with labels based on paper topics.
@@
 The Reddit dataset [52] consists of posts from different communities of the Reddit platform, where nodes represent posts, and edges connect posts commented on by the same user, forming a large interaction graph.
-Amazon Computer [68] is a co-purchase network, where nodes are products and edges indicate frequently co-purchased items within the computers category on Amazon.
+Amazon Computer [68] and Products [67] are co-purchase networks, where nodes are products and edges indicate frequently co-purchased items on Amazon.
 In contrast to previous dataset which are homophilous, Roman Empire [69] is an heterophilous dataset obtained from the homonymous Wikipedia page, where nodes represent words connected by syntactic relationships or adjacency in the sentence.
@@
 Only some of the nodes are labeled, either as licit (42,019 nodes) or illicit transaction (4,545 nodes).
-Summary statistics for the four graphs are reported in Table VIII.
-| Dataset | CoraFull | Arxiv | Amazon Computer | Roman Empire | Elliptic | |
-| # nodes | 19,793 | 169,343 | 227,853 | 13,752 | 22,662 | 203,769 |
-| # edges | 130,622 | 1,166,243 | 114,615,892 | 491,722 | 32,927 | 234,355 |
-| # classes | 70 | 40 | 40 | 10 | 18 | 2 |
-Appendix B Neighborhood expansion
-Neighborhood expansion has been identified as the main issue with the online setting for graphs.
-In Figure 2 of the main paper we showed the phenomenon, and we report here additional plots for all datasets and batch sizes.
-In particular, in Figures 6-8 we measure neighborhood expansion not only in terms of number of nodes in the -hop neighborhood, as done in Section VI of the main paper, but we additionally count the number of edges between hops, which gives us a good proxy of computational time.
-Interestingly, the number of edges increases more drastically than the number of nodes when incrementing the number of layers, further confirming the need for a strategy such as sampling to maintain the computational complexity bounded within the limits of the online learning setting.
-We do not present plots for the Roman Empire and Elliptic datasets, since their very low average degree (2.9 and 1.15 respectively) effectively eliminates the neighborhood expansion problem for a small number of layers.
-Appendix C Metrics
+Summary statistics for the four graphs are reported in Table X.
+| Dataset | CoraFull | Arxiv | Amazon Computer | Products | Roman Empire | Elliptic | |
+| # nodes | 19,793 | 169,343 | 227,853 | 13,752 | 2,449,028 | 22,662 | 203,769 |
+| # edges | 130,622 | 1,166,243 | 114,615,892 | 491,722 | 61,859,036 | 32,927 | 234,355 |
+| # classes | 70 | 40 | 40 | 10 | 46 | 18 | 2 |
+Appendix B Metrics
 Thanks to the construction of the node stream starting from the class-incremental setting (or the artificial definition of tasks for Elliptic), we can exploit two widely used metrics in CL: Average Performance (AP) and Average Forgetting (AF) [14].
@@
 Since the latter is highly unbalanced, on it we use the F1 score of the illicit class.
-To assess the performance of the model throughout the node stream, we also perform anytime evaluation, meaning that we evaluate the model on validation nodes after training on each mini-batch [28].
+More importantly, to assess the performance of the model throughout the node stream, we also perform anytime evaluation, meaning that we evaluate the model on validation nodes after training on each mini-batch [25].
 This allows us to capture the performance at any point in time, and observe also graphically how the model reacts to changes in data distribution.
@@
 This can be interpreted as an Area Under the Curve accuracy score [25].
-Appendix D Hyperparameters
+Appendix C Hyperparameters
 A standard grid search was performed to select training hyperparameters for the models used in all experiments.
@@
 No weight decay or dropout were used.
-Method specific hyperparameters are reported in Table IX, and specific details can be found in the original papers.
-In particular, the hyperparameters of regularization methods regulate the strenght of the regularization.
+Method specific hyperparameters are reported in Table XI, and specific details can be found in the original papers.
+In particular, the hyperparameters of regularization methods regulate the strength of the regularization.
 For LwF a new hyperparameter has been introduced to adapt to the online setting: the number of batches after which to update the teacher model.
@@
 | Method | Hyperparameter candidates |
-| A-GEM, ER, PDGNN | memory_proportion: {1,2,3} |
+| A-GEM, ER, PDGNN, LINEAR | memory_proportion: {1,2,3} |
 | SSM | neighbhors_budget: {(5,5), (10,10), (25,25)} |
@@
 | TWP | lambda_l: {}; lambda_t: {}; beta: {0.001, 0.01, 0.1} |
-Appendix E Extended results
-E-A Full results with larger batch size
-In Tables XI-XIII, we provide extended results with the same setting configurations as the one explored in Section VI of the main paper, yet with larger mini-batch size.
-Compared to size 10 for CoraFull, Amazon Computer and Roman Empire, and 50 for Arxiv, Reddit and Elliptic, we report here batch-sizes 50 and 250 respectively.
-| Method | AP% | AAPval% | AF% |
-| EWC | |||
-| LwF | |||
-| MAS | |||
-| TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
-| bare | |||
-| Joint | - | - |
-| Method | AP% | AAPval% | AF% |
-| EWC | |||
-| LwF | |||
-| MAS | |||
-| TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
-| bare | |||
-| Joint | - | - |
-| Method | AP% | AAPval% | AF% |
-| EWC | |||
-| LwF | |||
-| MAS | |||
-| TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
-| bare | |||
-| Joint | - | - |
-| Method | AP% | AAPval% | AF% |
-| EWC | |||
-| LwF | |||
-| MAS | |||
-| TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
-| bare | |||
-| Joint | - | - |
-| Method | AP% | AAPval% | AF% |
-| EWC | |||
-| LwF | |||
-| MAS | |||
-| TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
-| bare | |||
-| Joint | - | - |
-| Method | AP% | AAPval% | AF% |
-| EWC | |||
-| LwF | |||
-| MAS | |||
-| TWP | |||
-| ER | |||
-| A-GEM | |||
-| PDGNN | |||
-| SSM-ER | |||
-| SSM-A-GEM | |||
-| bare | |||
-| Joint | - | - |
-E-B Results with full neighborhood
-We report here empirical results obtained on mini-batches containing the full neighborhoods of newly presented nodes, in contrast with the main paper where neighborhood sampling was analyzed.
-Although the neighboring expansion problem is present, these results provide reference performance to be compared against the more realistic setting studied in the paper.
-For the Reddit dataset, due to the extremely high average degree, only one layer is used here for full neighborhood.
-Results are reported for CoraFull, Arxiv, Reddit and Amazon Computer datasets in Tables II-V, while they are not reported for Roman Empire and Elliptic as only a handful of nodes have a higher degree than 10 used for neighborhood sampling in the main experimentss of the paper.
-To be consistent with the use of full neighborhood information, the results reported for SSM are obtained by using the same number of seed nodes in the buffer, but with full neighborhood information as well.
-On the contrary, by using heavily sparsified subgraphs for the buffer in conjunction with full neighborhood for the node stream, we observed a degradation of performance, likely due to the different normalization weighting inside the GCN layer between the stream and the memory buffer.
-| Method | Batch size 10 | Batch size 50 | |||||
-| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
-| EWC | |||||||
-| LwF | |||||||
-| MAS | |||||||
-| TWP | |||||||
-| ER | |||||||
-| A-GEM | |||||||
-| SSM-ER | |||||||
-| SSM-A-GEM | |||||||
-| bare | |||||||
-| Joint | - | - | - | - | |||
-| Method | Batch size 50 | Batch size 250 | |||||
-| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
-| EWC | |||||||
-| LwF | |||||||
-| MAS | |||||||
-| TWP | |||||||
-| ER | |||||||
-| A-GEM | |||||||
-| SSM-ER | |||||||
-| SSM-A-GEM | |||||||
-| bare | |||||||
-| Joint | - | - | - | - | |||
-| Method | Batch size 50 | Batch size 250 | |||||
-| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
-| EWC | |||||||
-| LwF | |||||||
-| MAS | |||||||
-| TWP | |||||||
-| ER | |||||||
-| A-GEM | |||||||
-| SSM-ER | |||||||
-| SSM-A-GEM | |||||||
-| bare | |||||||
-| Joint | - | - | - | - | |||
-| Method | Batch size 10 | Batch size 50 | |||||
-| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
-| EWC | |||||||
-| LwF | |||||||
-| MAS | |||||||
-| TWP | |||||||
-| ER | |||||||
-| A-GEM | |||||||
-| SSM-ER | |||||||
-| SSM-A-GEM | |||||||
-| bare | |||||||
-| Joint | - | - | - | - | |||
-E-C Impact of buffer size
-For completeness of our results on the impact of buffer size on model performance, we show in Figures 10-11 plots similar to Figure 3 of the main body of the paper, showing both Average Performance and Average Anytime Performance achieved by replay methods with different buffer sizes, with both possible choices of mini-batch size as well.
-Appendix F Anytime evaluation plots
-We show here the plots with anytime evaluation, on all datasets and with the batch size used in the main experiments for the paper.
-In Figures 14-18 the lines indicate AP on validation nodes after training on each mini-batch, for all considered methods.
-In Figures 19-24, instead, we show the performance of each method in more detail, with a breakdown of accuracy by each task.
-Appendix G Ablation study on CoraFull
-In our experiments, the backbone model for all Continual Learning strategies was fixed as a 2-layer GCN with 256 hidden units, following Zhang et al.
-[46].
-Thus, we performed a small ablation study on the CoraFull dataset to assess the impact of the number of hidden units and number of GCN layers.
-This is conducted only with full neighborhood aggregation.
-G-A Number of hidden units
-Still maintaining a 2-layer GCN, we changed the number of hidden units, from the original 256 to alternatively 128 and 512.
-We report the results in Tables XX-XXI.
-We observe how overall performance are lower than with 256 hidden units, possibly as this number may have been validated in previous works.
-A smaller network seems to consistently damage performance metrics, while with 512 units the comparison is more mixed.
-| Method | Batch size 10 | Batch size 50 | |||||
-| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
-| EWC | |||||||
-| LwF | |||||||
-| MAS | |||||||
-| TWP | |||||||
-| ER | |||||||
-| A-GEM | |||||||
-| SSM-ER | |||||||
-| SSM-A-GEM | |||||||
-| bare | |||||||
-| Method | Batch size 10 | Batch size 50 | |||||
-| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
-| EWC | |||||||
-| LwF | |||||||
-| MAS | |||||||
-| TWP | |||||||
-| ER | |||||||
-| A-GEM | |||||||
-| SSM-ER | |||||||
-| SSM-A-GEM | |||||||
-| bare | |||||||
-G-B Number of GCN layers
-We considered here a different number of GCN layers compared to our main results, 1 and 3 specifically, as on CoraFull also the 3-hop neighborhood has a relatively limited expansion.
-Results are reported in Tables XXII-XXIII.
-We observe overall lower results specifically with 3 layers, while ER and A-GEM improve using only 1 GCN layer.
-Additionally, as a sort of “0-layer” baseline, we test the usage of a 2 layer MLP, equivalent to our main backbone GCN applied using the identity as adjacency matrix.
-This is the same as discarding graph topology information.
-For this reason, SSM is not used in this case.
-While most results are consequently greatly reduced, ER, GEM and EWC are able to maintain some performance even without using the graph topology.
-| Method | Batch size 10 | Batch size 50 | |||||
-| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
-| EWC | |||||||
-| LwF | |||||||
-| MAS | |||||||
-| TWP | |||||||
-| ER | |||||||
-| A-GEM | |||||||
-| SSM-ER | |||||||
-| SSM-A-GEM | |||||||
-| bare | |||||||
-| Method | Batch size 10 | Batch size 50 | |||||
-| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
-| EWC | |||||||
-| LwF | |||||||
-| MAS | |||||||
-| TWP | |||||||
-| ER | |||||||
-| A-GEM | |||||||
-| SSM-ER | |||||||
-| SSM-A-GEM | |||||||
-| bare | |||||||
-| Method | Batch size 10 | Batch size 50 | |||||
-| AP% | AAPval% | AF% | AP% | AAPval% | AF% | ||
-| EWC | |||||||
-| LwF | |||||||
-| MAS | |||||||
-| TWP | |||||||
-| ER | |||||||
-| A-GEM | |||||||
-| bare | |||||||
+Appendix D Ablation Study
+In the main experiments of the papers, some hyperparameter and design choices, such as batch sizes and the structure of the GCN model, are kept fixed.
+In this section we provide results to assess the impact of these choices, for three of the considered benchmarks: CoraFull, Amazon Computer and Arxiv (in class-incremental stream setting).
+For ease of reference, we report here their aggregated results from the tables in the main paper into Table XII.
+| CoraFull | Amazon Computer | Arxiv | |||||||
+| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
+| A-GEM | |||||||||
+| ER | |||||||||
+| EWC | |||||||||
+| LwF | |||||||||
+| MAS | |||||||||
+| PDGNN | |||||||||
+| SSM-A-GEM | |||||||||
+| SSM-ER | |||||||||
+| TWP | |||||||||
+| LINEAR | |||||||||
+| bare | |||||||||
+D-A Full results with larger batch size
+In Table XIII, we provide extended results with the same setting configurations as the one explored in Section VI of the main paper, yet with larger mini-batch size.
+Compared to size 10 for CoraFull and Amazon Computer, and 50 for Arxiv, here we use mini-batches of 50 and 250 nodes respectively.
+While on CoraFull and Arxiv the differences compared to the main setting are limited, on CoraFull we observe generally better performances (except for ER), with significantly higher results for regularization methods EWC and MAS.
+| CoraFull | Amazon Computer | Arxiv | |||||||
+| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
+| A-GEM | |||||||||
+| ER | |||||||||
+| EWC | |||||||||
+| LwF | |||||||||
+| MAS | |||||||||
+| PDGNN | |||||||||
+| SSM-A-GEM | |||||||||
+| SSM-ER | |||||||||
+| TWP | |||||||||
+| LINEAR | |||||||||
+| bare | |||||||||
+D-B Results with full neighborhood
+As one of the main requirements for the OCGL setting is to keep a constant computational footprint, in the experiments for the paper we adopt neighborhood sampling to tame the neighborhood expansion issue.
+Yet, we expect that this may come as a tradeoff with performance, since some information is thus discarded.
+To assess this, we report in Table XIV results obtained with mini-batches containing the full neighborhoods of newly presented nodes.
+We observe nonetheless how these results are in some cases lower than those with neighborhood sampling, suggesting that it could potentially act as a regularizer, and confirming the goodness of this efficient choice for the OCGL setting.
+| CoraFull | Amazon Computer | Arxiv | |||||||
+| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
+| A-GEM | |||||||||
+| ER | |||||||||
+| EWC | |||||||||
+| LwF | |||||||||
+| MAS | |||||||||
+| PDGNN | |||||||||
+| SSM-A-GEM | |||||||||
+| SSM-ER | |||||||||
+| TWP | |||||||||
+| LINEAR | |||||||||
+| bare | |||||||||
+D-C Number of hidden units
+Still maintaining a 2-layer GCN as in the main experiments, we changed here the number of hidden units, from the original 256 to alternatively 512 and 128.
+We report the results in Tables XV and XVI.
+LINEAR is not reported as it is a linear model.
+| CoraFull | Amazon Computer | Arxiv | |||||||
+| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
+| A-GEM | |||||||||
+| ER | |||||||||
+| EWC | |||||||||
+| LwF | |||||||||
+| MAS | |||||||||
+| PDGNN | |||||||||
+| SSM-A-GEM | |||||||||
+| SSM-ER | |||||||||
+| TWP | |||||||||
+| bare | |||||||||
+| CoraFull | Amazon Computer | Arxiv | |||||||
+| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
+| A-GEM | |||||||||
+| ER | |||||||||
+| EWC | |||||||||
+| LwF | |||||||||
+| MAS | |||||||||
+| PDGNN | |||||||||
+| SSM-A-GEM | |||||||||
+| SSM-ER | |||||||||
+| TWP | |||||||||
+| bare | |||||||||
+D-D Number of GCN layers
+We considered here a different number of GCN layers compared to the 2 our main results, 3 and 1 specifically, keeping the number of hidden units fixed to 256.
+Results are reported in Tables XVII and XVIII.
+LINEAR is not reported as it is a linear model.
+We observe overall lower results with 3 layers, in accordance with literature suggesting that deeper networks are more prone to forgetting [75].
+With only 1 layers results are mixed, with also some cases of improved performance over the 2-layer network.
+| CoraFull | Amazon Computer | Arxiv | |||||||
+| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
+| A-GEM | |||||||||
+| ER | |||||||||
+| EWC | |||||||||
+| LwF | |||||||||
+| MAS | |||||||||
+| PDGNN | |||||||||
+| SSM-A-GEM | |||||||||
+| SSM-ER | |||||||||
+| TWP | |||||||||
+| bare | |||||||||
+| CoraFull | Amazon Computer | Arxiv | |||||||
+| Method | AAP% | AP% | AF% | AAP% | AP% | AF% | AAP% | AP% | AF% |
+| A-GEM | |||||||||
+| ER | |||||||||
+| EWC | |||||||||
+| LwF | |||||||||
+| MAS | |||||||||
+| PDGNN | |||||||||
+| SSM-A-GEM | |||||||||
+| SSM-ER | |||||||||
+| TWP | |||||||||
+| bare | |||||||||
+Appendix E Performance visualization
+E-A Impact of buffer size
+For completeness of our results on the impact of buffer size on model performance for replay methods, for which Figure 5 of the main body of the paper shows AAP, we show here in Figure 6 the impact on AP.
+E-B Anytime evaluation plots
+We show here the plots with anytime evaluation, on all datasets and with the settings used in the main experiments for the paper.
+In Figures 8-14 the lines indicate AP on validation nodes after training on each mini-batch, for all considered methods.
+We highlight the boundaries between tasks, the threshold up to which hyperparameter selection is performed, and the upper bound of joint training up to the current task.
+We note how it is natural and expected that accuracy tends to decrease with the batch index, as new classes are introduced and the classification task gets increasingly complex.
+In general, we observe that while the performance of regularization methods tends to decrease smoothly, the performance of replay-based methods shows much higher variations.
+In Figures 15-22, instead, we show the performance of each method in more detail, with a breakdown of performance by each task.